### v3

# PyTomography: A Python Library for Medical Image Reconstruction Luke Polson1,2, Roberto Fedrigo1,2, Chenguang Li1,2, Maziar Sabouri1,2, Obed Dzikunu2,3, Shadab Ahamed1,2, Nicolas Karakatsanis4, Arman Rahmim1,2,3,5,∗, Carlos Uribe2,5,6,∗

Abstract
Background: There is an absence of open-source libraries in emission tomography that (i) use modern and popular backend code to encourage community contributions and (ii) offer support for the multitude of reconstruction algorithms available in recent literature, such as those that employ artificial intelligence.
Purpose: To create and evaluate a GPU-accelerated, open-source, and user-friendly image reconstruction library, designed to serve as a central platform for the development, validation, and deployment of various tomographic reconstruction algorithms.
Methods: PyTomography was developed using Python and inherits the GPU-accelerated functionality of PyTorch and parallelproj for fast computations. Its flexible and modular design decouples the system matrix from reconstruction algorithms, simplifying the process of integrating new imaging modalities using various python tools.
Results: Example use cases demonstrate the software capabilities in parallel hole SPECT and listmode PET imaging. Use of the OSEM, BSREM, KEM, and DIPRecon reconstruction algorithms are explored. System matrices incorporating breathing-motion correction are explored in SPECT imaging. PyTomography is validated against alternative vendor/open source software for OSEM based reconstruction.
Conclusions: We have developed and publicly shared PyTomography, a highly optimized and user-friendly software for medical image reconstruction, with a class hierarchy that fosters the development of novel imaging applications.
I. Introduction
Medical imaging forms a cornerstone in modern healthcare by providing visual and quantitative information about internal body structures and functions. It enables early detection1, accurate diagnosis, and precise treatment planning for a wide range of medical conditions2, 3, 4, contributing to improved patient outcomes and enhanced medical decision-making. Reconstruction algorithms are routinely used to generate 3D images that can be used for both research and clinical decision making 5.
While the development and validation of reconstruction techniques remains an active field of research6, 7, 8, 9, it can often be difficult to share and disseminate one’s findings. Additionally, while manufacturers of tomographic imaging equipment such as Single Photon Emission Computed Tomography (SPECT) and Positron Emission Tomography (PET) cameras provide their own internal reconstruction software, users cannot access all implementation details; this limits reproducibility in scientific studies. These issues could be solved by moving to an open source paradigm for medical image reconstruction. Here, the imaging community could openly collaborate in the development of reconstruction techniques. Adoption of a singular open-source framework, as opposed to private frameworks developed independently by various research groups, would strengthen the ability to develop, share, and compare findings. This could ultimately help accelerate the development and translation of new diagnostic imaging capabilities into clinical practice.
A few tools have already been developed to try and address some of these issues. STIR 10 and CASToR 11 are two C/C++ based open source and collaborative image reconstruction frameworks that support a variety of reconstruction algorithms for SPECT and PET. NiftyPET 12 is a python library which employs GPU-accelerated code written in CUDA C for PET reconstruction. QSPECT13 and TIRIUS support use of the ordered subset expectation maximum (OSEM) algorithm for SPECT and PET reconstruction respectively. Other libraries, such as TomoPy 14 and the ASTRA toolbox 15, support a variety of GPU-accelerated image tomographic reconstruction algorithms, although they lack extensive SPECT/PET system modeling. Web links to all these tools, as well as a list of other available tools for image reconstruction, can be found on the NMMItools webpage of the Society of Nuclear Medicine and Molecular Imaging (SNMMI). The unique purpose of this research is to build upon recent efforts in GPU-accelerated reconstruction16, 17, 18, 19 and AI-based reconstruction7, 8, 9 to establish an open source medical image reconstruction programming framework written in a popular programming language that facilitates collaboration. Valuable past efforts have focused on CUDA programming to accelerate reconstruction times in SPECT16 and PET 16, 17 reconstruction through the use of multi-threading. However, potential contributors to an open-source framework may be unfamiliar with CUDA; any tool that prioritizes community contributions should employ familiar backend code. In Python, the library NumPy is widely known, but it lacks direct GPU acceleration support. PyTorch, however, emerges as a suitable alternative since it maintains a similar interface to NumPy while also facilitating GPU acceleration and direct interface with neural networks. Some potential contributors may wish to provide contributions using GPU-accelerated functionality offered by other libraries, such as CuPy; an optimal framework should thus also permit the ability to incorporate alternative implementations for particular tasks.
To this end, the python library PyTomography has been developed with the following priorities: (i) implementation of standard and traditional imaging modalities and reconstruction algorithms, (ii) implementation and dissemination of recent research developments, such as the deep image prior8, and (iii) encouragement of community involvement via extensive documentation and user-friendly tutorials. While the present focus of PyTomography is the development and validation of reconstruction algorithms for SPECT/PET, implementation of other imaging modalities and reconstruction algorithms can be added using the building blocks provided.
Overall, the aim of this study is to introduce the community to the PyTomography package, demonstrate its current capabilities and flexibility, and to encourage community involvement in the continuation of its development. To demonstrate the versatility of the software, example applications will be explored in (i) SPECT reconstruction using PyTorch based rotate+sum projectors, and (ii) listmode (LM) time of flight (TOF) PET reconstruction using Joseph’s projectors from the python library parallelproj19. In what follows, we elaborate on our methods and results for PyTomography, which is made publicly available to download; relevant links can be found in section VI..
II. Materials and Methods
II.A. Mathematical Description
A brief overview of the mathematical paradigm of medical imaging is presented in order to motivate the software architecture. The mathematical conventions used in this paper are as follows. Vectors are represented by lower case letters, while linear operators are represented by upper case letters. The product of two vectors $vw$ and division of two vectors $v/w$ are inferred to be point-wise operations. Components of a vector are given by $v_{i}$, and components of linear operators are given by $A_{ij}$.
II.A.1. Projectors
In the paradigm of tomographic medical imaging, there are two main vector spaces: the “projection” space $\mathcal{V}$ consisting of measured projection data $g$ and the “object” space $\mathcal{U}$ consisting of 3D objects $f$. Since measured and reconstructed data is digitized, $\mathcal{U}$ and $\mathcal{V}$ are assumed here to be finite dimensional vector spaces. In SPECT imaging, for example, $\mathcal{V}$ consists of a set of 2D pixelated images acquired by counting photons emitted by radiopharmaceuticals, and $\mathcal{U}$ represents the voxelized spatial distribution of radioactivity concentration in the object being imaged.
Certain imaging systems can be characterized by a system matrix $H$, which corresponds to a linear operator mapping from object space to projection space. Its components $H_{ij}$ represent a contribution of voxel $j$ to detector element $i$. In SPECT/PET, analytical modeling of the data expectation $\bar{g}$ can be achieved via $\bar{g}=Hf+s$ where $s$ represents scatters and randoms. In CT, analytical modeling is achieved via $\bar{g}=be^{-Hf}$ where $b$ denotes the rate of emissions from the transmission source to a particular measurement. The operation of $H$ on an object vector is known as forward projection (FP), while the operation of $H^{T}$ on a projection vector is known as back projection (BP).
Efficient implementation of FP and BP depends on the particular imaging modality; this typically involves decomposing $H$ into a sequence of less computationally burdensome operations. In general, FP can be decomposed into the following sequence: (i) linear endomorphisms mapping from $\mathcal{U}\to\mathcal{U}$, (ii) operators mappings from $\mathcal{U}\to\mathcal{V}$, and (iii) linear endomorphisms mapping from $\mathcal{V}\to\mathcal{V}$ given by $\{B_{k}\}$. BP is given by the transpose of this operation sequence. In PyTomography, operators which map to the same vector space are defined as transforms, while operators that map between different vector spaces are defined as projectors. Example implementations available in PyTomography are listed below:
  • Parallel Collimator SPECT: Modeling of primary detections in this imaging modality can be approximated using the “rotate+sum” technique16, whereby (i) consists of rotations and attenuation/collimator detector response modeling, (ii) consists of summation along an axis, and (iii) consists of adjustments for a camera with a finite field of view. System modeling is approximated using the following system matrix:
    $H=\sum_{n}v_{n}\otimes\left(\mathcal{T}_{\text{fov}}^{(n)}~{}\mathcal{P}_{x}~{}\mathcal{T}_{\text{cdr}}^{(n)}\mathcal{T}_{\text{att}}^{(n)}~{}\mathcal{T}_{\text{trans}}^{(n)}~{}\mathcal{T}_{\text{rot}}^{(n)}\right)$ (1)
    where $n$ designates the projection angle, $\mathcal{T}_{\text{rot}}^{(n)}$ (transform) implements a rotation, $\mathcal{T}_{\text{trans}}^{(n)}$ (transform) implements a translation that accounts for non-central projections, $\mathcal{T}_{\text{att}}^{(n)}$ (transform) implements the adjustment for photon attenuation, $\mathcal{T}_{\text{cdr}}^{(n)}$ (transform) implements depth-dependent collimator blurring, $\mathcal{P}_{x}$ (projector) is summation along the $+x$ axis, and $\mathcal{T}_{\text{fov}}^{(n)}$ (transform) applies a mask in projection space that accounts for a finite field of view, $v_{n}$ is a unit vector of length $N$ with 1 in component $n$, and $\otimes$ is the outer product.
  • LM PET: Modeling in this modality can be approximated when (i) consists of object-space resolution modeling, (ii) consists of ray tracing between all detected lines of response (LOR), and (iii) consists of normalization/attenuation correction for each LOR. In this case, the system matrix can be written as
    $H=\mathcal{T}_{\eta}~{}\mathcal{T}_{\omega}~{}\mathcal{P}_{\Sigma}~{}\mathcal{T}_{rr}$ (2)
    where $\mathcal{T}_{rr}$ (transform) is used for object-space resolution modeling, $\mathcal{P}_{\Sigma}$ (projector) is used to compute line integrals, $\mathcal{T}_{\eta}$ (transform) adjusts for detector normalization, and $\mathcal{T}_{\omega}$ (transform) adjusts for attenuation effects. Due to cancellation in certain reconstruction algorithms, $\mathcal{T}_{\eta}$ and $\mathcal{T}_{\omega}$ need not always be included (though they are required for computing the sensitivity factor).
PyTomography can be used to construct projectors that extend projection space; such functionality is necessary, for example, when modeling multiple photopeaks in SPECT or when using motion correction. In such situations, the extended projection space has dimensionality $(N,...)$ where ... is the original dimensionality of projection space, and $N$ represents the number of photopeaks or time gates during acquisition. Mathematically, these projectors can be written as
$H^{\prime}=\sum_{n}v_{n}\otimes B_{n}H_{n}A_{n}$ (3)
where $\left\{A_{n}\right\}$ are object-to-object space transforms, $\left\{H_{n}\right\}$ are a sequence of system matrices, $\left\{B_{n}\right\}$ are a sequence of projection-to-projection space transforms, $v_{n}$ is a basis vector of length $N$ with a value of $1$ in component $n$ and $\otimes$ is an outer product. This can be utilized in the following examples:
  • Periodic Motion Correction: $A_{n}$ represent spatial transformation operators corresponding to time gate $n$, $H_{n}$ represent system matrices for each motion transform (attenuation modeling may depend on object deformation), and $B_{n}=1$. This maps to an extended projection space $(N,...)$ where $N$ is the number of different time gates (e.g. during a breathing cycle).
  • Multiple SPECT Photopeaks: In this scenario, $A_{n}$ are scaling factors that adjust for detector sensitivities and branching ratios, $B_{n}$ are identity mappings, and $H_{n}$ represents the system matrix corresponding to a particular photopeak. The extended projection space has dimensions $(N,...)$ where $N$ is the number of different photopeaks.
II.A.2. Likelihoods
A likelihood function $L(g|\bar{g})$ provides the likelihood of observing data $g$ given expectation $\bar{g}$. Since $\bar{g}$ is approximated via analytical models depending on $f$ and $H$, it follows that $L(g|\bar{g})=L(g|f,H,...)$, where the ellipsis “...” contains additional parameters such as a scatter estimate $s$; the ellipsis will assumed implicitly for the remainder of this paper. Many iterative reconstruction algorithms seek to maximize the likelihood function via gradient ascent using the gradient $\nabla_{f}L(g|f,H)$. Likelihood functions for different modalities are shown in Table 1.
Table 1: Likelihoods and corresponding gradients for various medical imaging modalities; $s$ represents a scatter estimate used in SPECT/PET while $b$ represents a scaling factor used in CT.
Likelihood $L(g|f,H)$ $\nabla_{f}L(g|f,H)$
SPECT/PET $\sum_{i}\left(g_{i}\log\left[Hf+s\right]_{i}-\left[Hf+s\right]_{i}\right)$ $H^{T}\left(\frac{g}{Hf+s}\right)-H^{T}1$
CT $\sum_{i}\left(g_{i}\log\left[be^{-Hf}\right]_{i}-\left[be^{-Hf}\right]_{i}\right)$ $H^{T}(be^{-Hf}-g)$
Weighted MSE $-\frac{1}{2}\sum_{i}\left[W\left(g-Hf\right)^{2}\right]_{i}$ $H^{T}W(g-Hf)$
II.A.3. Reconstruction Algorithms
An image reconstruction algorithm $A$ uses measured data $g$ (i.e. projections) to estimate a corresponding object $f$ that would produce projections $g$ given a system matrix $H$. Statistical algorithms find the image estimate $f$ which maximizes a likelihood function $L(g|H,f)$ characterizing the imaging system. This can be expressed as
$f=A(L,...)$ (4)
where the ellipsis “...”includes all additional hyperparameters required for the algorithm; the ellipsis will be assumed implicit for the remainder of this paper. The main algorithms in PyTomography are preconditioned gradient ascent (PGA) algorithms; these algorithms aim to maximimize the likelihood via iterative updates that take the form
$f^{n+1}=f^{n}+C^{n}(f^{n},H_{n})\left[\nabla_{f}L(g^{n}|f^{n},H_{n})-\beta_{n}\nabla_{f}V(f^{n})\right]$ (5)
where $n$ is the subiteration index, $f^{n}$ is the object estimate of subiteration $n$, $g^{n}$ specifies a subset of data used in subiteration $n$, $H_{n}$ is a system matrix projecting only to the current subset, $C^{n}(f^{n})$ is a preconditioner factor that depends on the object estimate, $V(f^{n})$ is a prior function used to regularize the algorithm, and $\beta_{n}=c_{n}\beta$ is a scaling factor for the prior function (scaled by the fraction of subset data $c_{n}$). These algorithms partition the data $g$ into $M$ approximately equal subsets; a single iteration of an algorithm is defined as the number of subiterations where all subsets have been used once. Various PGA algorithms available in PyTomography are shown in Table 2; they are characterized by their preconditioner factor. Additional reconstruction techniques can be developed using these algorithms. The kernel expectation maximum (KEM), for example, corresponds to (i) reconstructing a coefficient image $\alpha$ using $\alpha=A_{\text{OSEM}}(g,HK)$ where $K$ is a kernel transform and (ii) obtaining the image estimate via $f=K\alpha$.
Table 2: Sample algorithms available in PyTomography: shown are ordered subset expectation maximum (OSEM), ordered subset maximum a posteriori one step late (OSMAPOSL), block sequential regualized expectation maximum (BSREM), and kernelized expectation maximum (KEM).
Algorithm Preconditioner factor $C^{n}(f^{n},H)$
OSEM20 $f^{n}/\left(H_{n}^{T}1\right)$ and $V(f^{n})=0$
OSMAPOSL21 $f^{n}/\left(H_{n}^{T}1+\nabla_{f}V(f^{n})\right)$
BSREM22 $f^{n}\alpha(n)/\left(\omega_{n}H^{T}1\right)$
Certain algorithm names in the literature correspond to specific realizations of the generic algorithms in Table 2 with particular likelihoods in Table 1. The OS-SART algorithm, for example, corresponds to the use of OSEM with the weighted MSE likelihood (with weights $W=1/(H1)$).
Prior functions $V(f)$ used in reconstruction often take the form
$V(f)=\sum_{r}\sum_{s}w_{rs}\phi(f_{r},f_{s})$ (6)
where $w_{rs}$ provides a weighting factor between the locations voxel $r$ and voxel $s$ and $\phi(f_{r},f_{s})$ characterizes a weighting factor between the observed image values at voxel $r$ and voxel $s$. When $w_{rs}$ is non-zero only for neighbouring voxels, the prior is known as a nearest neighbour prior; various nearest neighbour priors implemented in PyTomography are shown in Table 3.
Table 3: Sample nearest neighbour functions available in PyTomography; these priors are characterized by their image weighting function $\phi(f_{r},f_{s})$.
Prior $\phi(f_{r},f_{s})$
Quadratic $\left((f_{r}-f_{s})/\delta\right)^{2}/4$
Log-cosh $\tanh((f_{r}-f_{s})/\delta)$
Relative difference penalty (RDP) 24 $\left((f_{r}-f_{s})^{2}\right)/\left(f_{r}+f_{s}+\gamma|f_{r}-f_{s}|\right)$
While the weighting factor $w_{rs}$ is typically proportional to inverse Euclidean distance, it is also possible to allow it to depend on an additional anatomical image $\chi$ whereby $w_{rs}=w_{rs}(\chi_{r},\chi_{s})$ Since anatomical information is used, this is commonly referred to as an anatomical prior (AP); variations of this technique have recently become popular in clinical practice due to enhanced lesion quantitation and detectability in bone SPECT/CT6. In this paper, use of an AP refers to using only the 8 most similar neighbours based on absolute differences of HU values in a CT scan or attenuation map.
PyTomography also offers support for the Deep Image Prior Reconstruction (DIPRecon) algorithm. Details regarding the implementation of this algorithm can be found in Algorithm 2 of Gong et. al8. The purpose of this reconstruction technique is to use anatomical images (CT/MR) to enhance functional images (SPECT/PET) through a sequence of neural network overfitting and mixing with OSEM updates. This reconstruction algorithm requires the use of an anatomical image $z$ and a neural network architecture $\mathcal{N}(\cdot;\theta)$ that can be trained to predict a reconstructed functional image $f$ given $z$ as input. This paper explores the use of DIPRecon on an ultra-high resolution PET/MR brain phantom.
II.B. Software Architecture
The key points of the last section are that (i) medical imaging consists of many different modalities that each require unique implementations for system modeling and (ii) tomographic reconstruction algorithms are modality agnostic. PyTomography has been carefully designed to permit arbitrary implementations of system modeling that can interface with the available reconstruction algorithms. The components of PyTomography, designed to facilitate the architecture described, are listed below.
  1. 1.
    Data. The standard data type used to represent images and raw projection data is the tensor class of PyTorch.
  2. 2.
    Metadata. Metadata classes contain all auxiliary information needed for image reconstruction pertaining to a particular imaging system, such as voxel size, spatial dimensions, and information pertaining to the raw data of a particular modality.
  3. 3.
    Transforms. Transform classes are used as utilities to model various features of the imaging process, such as attenuation and image blurring in SPECT. They are commonly used as components in a system matrix. Specifically, transforms are defined by the mathematical property that they are linear endomorphisms.
  4. 4.
    Projectors. Unlike transforms, projectors are mathematical operators that map between different vector spaces. The system matrix, which maps from object space to projection space and represents a model of the imaging system, is an example of a projector. While the input/output datatype to projectors is the PyTorch tensor class, internal code can utilize functionality from other libraries.
  5. 5.
    Likelihoods. Likelihoods specify the likelihood function that characterizes the acquired data. Likelihoods $L(g|H,f)$ are initialized using a system matrix $H$, projection data $g$, and any additional dependencies; once initialized they act as functions of an image estimate $f$.
  6. 6.
    Algorithms. Algorithms are used for image reconstruction, and are designed to interface with all imaging modalities. Statistical iterative algorithms require as input a likelihood instance, and may take as optional arguments prior functions and callbacks.
A visual representation of the software library along with sample code representing a SPECT reconstruction use case is shown in Figure 1. Metadata is first obtained via the raw vendor data stored in the path represented by the variable “file_NM”. The metadata is used as input to a “SPECTSystemMatrix” class, which also takes as input attenuation, PSF, and field of view cutoff transforms (creation of transforms not shown). The generated system matrix is used to build a likelihood instance, which also takes in as input the photopeak (creation not shown) and scatter as an additive term (creation not shown). The likelihood is then used to initialize an OSEM class for reconstruction; a reconstructed object is subsequently obtained using 4 iterations and 8 subset.
Figure 1: Basic overview of the workflow of PyTomography (top) with code representing a specific SPECT reconstruction use case (bottom). Sample classes and implementations are shown; these do not comprise the full set of classes available in the software.
II.C. Examples
Multiple examples are explored to highlight the capabilities of PyTomography. At present, reconstruction of parallel collimator SPECT and LM PET data is demonstrated, but it should be emphasized that these are purely intended as example use cases for particular imaging modalities. All computations were performed using a Microsoft Azure virtual machine (Standard NC6s v3) with a 6 CPUs (Intel(R) Xeon(R) CPU E5-2690 v4 @ 2.60GHz), 112 GB of RAM, and a TeslaV100 GPU. Python scripts and notebooks used to obtain these results can be found at https://github.com/qurit/PyTomography_paper_code.
II.C.1. Monte Carlo SPECT
Two use cases are explored for reconstruction of Monte Carlo generated SPECT image data from the SIMIND Monte Carlo simulation program26. The system matrices used in each case correspond to equation 1, with each employing attenuation modeling and PSF modeling. The triple energy window technique27 is used for scatter correction.
[…]
Table 4: Reconstruction parameters used for SPECT reconstruction in this paper. Neighbourhood size corresponds to the kernel size in KEM, and prior neighbourhood for BSREM. Nearest neighbours were obtained from attenuation maps.
Algorithm OSEM BSREM KEM
Subsets 8 8 8
Neighbourhood Size - $3\times 3\times 3$ $5\times 5\times 5$
$k$ Nearest Neighbours - 8 40
Kernel - - Eq. 5 of Vuohijoki et. al. 6
Prior - RDP-AP -
[…]
II.C.2. Clinical SPECT
This example highlights SPECT reconstruction on publicly available data from the Deep Blue data repository31 that was used for the SNMMI 177Lu dosimetry challenge32, 33. Patients received radiopharmaceutical therapy with 177Lu-DOTATATE for neuroendocrine tumours (4 cycles of 7.4 GBq/cycle administered every 8 weeks); scans were taken at 4 different time points in the week following a therapeutic dose. Images consisted of 120 projections of shape $256\times 256$ with resolution $2.4~{}\text{mm}\times 2.4~{}\text{mm}$. The system matrix employed attenuation, PSF, and finite FOV modeling. Reconstructions using OSEM, BSREM, and KEM were performed for 50 iterations each; the TEW technique was used to correct for scatter in all cases.
II.C.3. Monte Carlo PET
This example demonstrates (i) the capabilities of LM TOF PET reconstruction on data generated by the GATE Monte Carlo simulation program and (ii) the DIP Recon technique. The simulated scanner consisted of $8$ rings of 338 mm radius with 32.25 mm ring spacing, and a TOF coincidence resolution of 550 ps. For simplicity, data acquisition in the simulation ignored dead-time effects and only considered true primary coincidence events; coincidences due to phantom-scatter and randoms were ignored. The digital phantom used in simulation corresponded to an ultra-high resolution 18F-FDG PET/ T1-weighted MRI brain image34; the full simulated dataset was representative of a 3 min scan and consisted of $556$ million coincidence events. Data was reconstructed using a $128\times 128\times 96$ matrix size with voxel spacing of 222~{}2mm. The normalization $\eta$ was obtained via simulation of a thin cylindrical shell of radius 318318318~{}318mm that extended along the axial FOV of the scanner; normalization computation is available in PyTomography for basic polygonal PET geometries. Reconstruction was performed with normalization correction in 3 different ways:
  1. 1.
    The full dataset (556 million events) was reconstructed using OSEM (4it, 20ss) using the EM algorithm
  2. 2.
    6.3% of the data set (35 million events) was reconstructed using OSEM (80it, 1ss) using the EM algorithm, representative of a low-count scenario.
  3. 3.
    6.3% of the dataset (35 million events) was reconstructed using the DIPRecon algorithm. The prior image $z$ used was the corresponding T1-weighted MRI image; the initial network prediction $\mathcal{N}(z|\theta_{0})$ was obtained via fitting to the previous low count reconstruction. The neural network was a UNet35 with similar architecture to the original DIP paper 8; more information on the network structure and reconstruction hyperparameters can be found in the publicly available code.
III. Results
The time required for reconstruction in the first SPECT use case was 412.8 s (3.44 s/iteration) for OSEM, 434.8 s (3.62 s/iteration) for BSREM, and 601.8 s (5.02 s/iteration) for KEM; sample reconstructions with corresponding RC curves are shown for each algorithm in Figure 2. BSREM clearly produced the least noisy images, and had the best RCs for the lungs, liver, kidneys, and bladder. KEM produced an image estimate with an intermediate level of noise, and had the best recovery coefficient for the salivary glands. OSEM yielded the noisiest images, and had the worst recovery coefficients in all regions except the salivary glands. Despite the high performance of BSREM, it was unable to accurately predict the regions of low uptake within the liver visible in the ground truth.
Figure 2: Reconstructions of full body 177Lu-PSMA-617 Monte Carlo projection data; ground truth phantoms were obtained using the XCAT software and SPECT simulation was achieved using the SIMIND Monte Carlo program. Left: Central coronal slices corresponding to truth, OSEM (120it/8ss), BSREM (120it/8ss), and KEM (120it/8ss). Right: Recovery coefficient vs. iteration number for sample organs; line widths are representative of error estimates based on reconstructions of 5 independent noise realizations.
Sample reconstructions and RC curves for the SPECT motion correction use case are shown in Figure 3. The lesions appear sharper when motion correction is used; motion correction applied during reconstruction via Equation 3 produced lesions of higher uptake than when correction was applied post-reconstruction. This is quantitatively observed in the RC curves of Figure 3, where motion correction applied during reconstruction tends to produce the highest RCs throughout all iterations. After 100 iterations, the RCs were larger when MC was used during reconstruction than when it was used post reconstruction ($p<10^{-6}$).
Figure 3: Reconstructions of full body 177Lu-PSMA-617 Monte Carlo projection data with breathing motion simulated; ground truth phantoms were obtained using the XCAT software and SPECT simulation was achieved using the SIMIND Monte Carlo program. Reconstruction was performed with OSEM (8ss) up to 100 iterations using attenuation/PSF modeling and TEW scatter correction; images shown correspond to reconstruction after 20 iterations. From left to right: ground truth phantom, reconstructed image without MC, reconstructed image with MC applied after reconstruction, reconstructed image with MC applied during reconstruction, and lesion recovery coefficients as a function of iteration number. Green boxes are drawn around the lesion in the liver.
The time required to reconstruct the clinical SPECT patient data was 464.7s (9.29 s/iteration) for OSEM, 453.1s (9.06 s/iteration) for BSREM, and 1825s (36.50 s/iteration) for KEM. Central coronal slices are shown in Figure 4; the patient had a region of high uptake value in the liver, indicative of a liver lesion. BSREM and KEM appear to confine activity within the boundary of the kidneys, with BSREM exhibiting slightly more of this behaviour. In addition, OSEM produced the most observed noise in the liver. A common artifact of both BSREM and KEM, however, is the noise in the liver lesion; use of anatomical nearest neighbours may cause the lesion activity noise profile to resemble the CT noise profile. The mean predicted counts in the kidneys were 1.98 (OSEM), 2.07 (BSREM), and 2.02 (KEM); these relative proportions are consistent with the RCs of Figure 2, suggesting that BSREM may yield the most accurate recovery coefficients for the kidneys.
Figure 4: Coronal slices of activity concentration (colored) and Hounsfield Unit values (greyscale) reconstructed using OSEM (50it/8ss), BSREM (50it/8ss), and KEM (50it/8ss). Data corresponds to the first time point of patient 4 from the “Lu-177 DOTATATE Anonymized Patient Datasets: Multi-Time Point Lu-177 SPECT/CT Scans” dataset31.
Reconstructing the LM TOF PET GATE output data required multiple steps. While initial processing of the ROOT listmode data is time consuming, PyTomography can save this data in an accessible format for reuse.
  • Opening the ROOT list mode data ($\sim$ 50 GB) of the normalization scan and computation of the normalization factor $\eta$ (approximately 50 minutes). When the normalization factor is saved, it requires approximately 5 s to read.
  • Opening the ROOT list mode data ($\sim$ 74 GB) of the phantom scan (approximately 1 hour). When saved in an accessible format, reading requires approximately 5 s.
  • Initial configuration of the system matrix ($7.8$ s).
  • Reconstruction of the PET listmode data. The high count scenario using OSEM (4it 20ss) required 185.6 s (46.4 s per iteration), the low count scenario using OSEM (80it 1ss) required 201.0 (2.5 s per iteration), and the low count scenario using DIPRecon required 3135.6 s (31.3 s per iteration)
Corresponding phantom reconstructions are shown in Figure 5; bias and noise in grey and white matter is shown as a function of iteration number in Figure 6. Due to finite resolution of the imaging system causing activity “leakage”, activity quantification in the grey matter region was consistently underestimated while activity quantification in the white matter region was consistently overestimated. The DIPRecon on the low count data had similar noise levels in grey matter to OSEM on the high count data, demonstrated both qualitatively and quantitatively in each Figures 5 and 6 respectively.
Figure 5: PET reconstruction of GATE simulated data corresponding to a ultra-high resolution 18F-FDG PET/ T1-weighted MRI brain image34. From top to bottom: axial, coronal, and sagittal slices. From left to right: ground truth MRI (non contrast adjusted), ground truth PET uptake, high count OSEM reconstruction (20it/4ss), low count OSEM reconstruction (80it/1ss), and low count DIP reconstruction.
Figure 6: Variation in absolute bias and standard deviation throughout iterative reconstruction corresponding to reconstructed images of Figure 5. Subplots correspond to the grey matter region (left) and white matter region (right). For the low count rate scenario, points are plotted every 10 iterations. For the low count rate scenario with OSEM, the first point corresponds to the 10th iteration. The bias was negative in grey matter and positive in white matter.
IV. Discussion
This technical report presented the software architecture of PyTomography, and showed use cases on both SPECT and PET data. The short times required for reconstruction using this tool can enable researchers to perform extensive studies on phantom and patient data that may include multiple patients, noise realizations, and reconstruction algorithms. Validation of PyTomography against alternative reconstruction libraries is shown in the appendix.
Example applications in 177Lu-SPECT imaging, which are relevant for currently FDA approved radiopharmaceutical therapies, explored the use of the OSEM, BSREM, and KEM, reconstruction algorithms. It should be emphasized that the purpose of this technical report was not to rank the different algorithms, but rather to demonstrate the capabilities of the software library. While some evidence was shown to favour BSREM over OSEM and KEM for kidney and liver dosimetry based on better recovery coefficients (Figure 2), a truly comprehensive study that aims to quantitatively compare algorithms should involve (i) multiple XCAT activity and anatomical configurations, (ii) inclusion of patient motion, (iii) effect of misalignment between SPECT and CT images, (iv) different hyperparameters in each reconstruction algorithm and (iv) should seek to use real patient data to validate any phantom studies. In addition, recovery coefficients alone are not sufficient to rank algorithms. Studies should also consider source-to-background ratios as well as error analysis based on intra- and inter-patient variabilities, and clinical task performance. While the presented software framework enables such studies to be performed, it is out of scope for this publication.
While the RDP-AP prior was featured heavily in this study, the framework of PyTomography is designed to foster the development and validation of other novel prior functions as well. In addition, while KEM was featured using anatomically-based basis functions, the library permits use of any external image for construction of the basis functions. Different priors and basis functions may also have significant implications for (i) quantitative PET/SPECT imaging in dosimetry, and (ii) qualitative observer based studies in lesion detection and classification.
The motion correction example considered a simplified case of 4 discrete phases where motion correction transforms were obtained using ground truth attenuation maps. In practice, however, breathing motion is continuous, and motion correction transforms will need to be derived using alternative source of data. The scope of this technical report is merely to demonstrate that PyTomography contains the functionality for incorporating motion correction.
Example applications in 18F-FDG PET imaging demonstrated the capabilities of LM reconstruction using the OSEM and DIPRecon algorithms. As demonstrated qualitatively in Figure 5 and quantitatively in Figure 6, reconstruction of the low count data using DIPRecon was able to reduce noise in the grey matter regions while preserving cortical structures. Once again, the purpose of this paper is merely to demonstrate the capabilities of PET LM reconstruction and use of the DIPRecon algorithm provided in the software library; extensive research on the validity of the DIPRecon method should seek to explore, for example, the effects of different neural network architectures, hyperparameters, and misalignment between the PET/MRI images.
While this paper has demonstrated implementations for parallel collimator SPECT and LM TOF PET, there exist many other modalities not presently included in the software, including (but not limited to) diverging/converging/pinhole SPECT, sinogram-based PET reconstruction, and various forms of computed tomography (CT) and magnetic resonance (MR) imaging. As displayed in Figure 1, projectors for different imaging modalities can be implemented using a variety of python based tools: it is hoped that the structure and flexibility of PyTomography will encourage contributions from various research groups interested in particular imaging systems. Since the library is open source, newly developed functionality can be easily tested and verified by many independent research groups with their own data. The immediate goals for future development in PyTomography at the time of publication include:
  1. 1.
    Fast PSF modeling for high energy SPECT isotopes, such as 131I and 225Ac.
  2. 2.
    Incorporation of scatter and random estimation in PET reconstruction.
  3. 3.
    Sinogram-based PET reconstructions.
  4. 4.
    Development of a 3D Slicer36 extension.
In summary, we hope that we have made a compelling case for the benefit that this tool can bring to the medical imaging community. It can be used as a framework for studies that require fast reconstruction times, such as those that explore a wide range of reconstruction hyperparameters. It is well validated and may be used for accurate and reproducible quantitative imaging in both PET and SPECT; we particularly envision applications in the field of theranostics. A variety of developments will assist PyTomography in reaching the full potential of precision medicine. We thus highly encourage and welcome open-source contributions to assist in the items listed above; those who wish to contribute are encouraged to read the developers’ guide on the documentation website.
V. Conclusion
This work describes the python library PyTomography and highlights specific use cases in SPECT and PET imaging. The software architecture facilitates the development of different imaging modalities that can all interface with the same reconstruction algorithms. The purpose of PyTomography is to create a well-documented and computationally-efficient library for medical image reconstruction, where novel system modeling and reconstruction techniques are implemented, shared, and explored by experts in the community.
VI. Appendix
VI.A. Relevant Links
Links to the various webpages of PyTomography are contained in Table 5: GitHub contains the source code, Readthedocs contains all relevant documentation, PyPI is the host of the built source files, and Discourse is a community forum for users and contributors.
Table 5: Relevant Webpages
Website Link
Github https://github.com/qurit/PyTomography
Readthedocs https://pytomography.readthedocs.io/en/latest/
PyPI https://pypi.org/project/pytomography/
Discourse https://pytomography.discourse.group/
VI.B. Validation of PyTomography
This section contains two examples used to validate PyTomography against other reconstruction software:
  1. 1.
    Monte Carlo SPECT Validation With STIR: Radioactivity concentrations in a digital phantom were selected to correspond to a 24-hour post injection prostate cancer patient (1700 MBq total activity) imaged via the prostate-specific membrane antigen (PSMA)-targeting radiopharmaceutical, 177Lu-PSMA-617. Simulated data was reconstructed with OSEM (2 iterations, 8 subsets) using (i) STIR and (ii) PyTomography. Coronal slices and 1D profiles of reconstructions are shown in Figure 7 and are nearly identical. The image noise, quantified by the standard deviation of predicted counts in the liver, was 0.789 counts (STIR) and 0.793 counts (PyTomography). The mean predicted counts in the kidneys were 1.764 (STIR) and 1.761 (PyTomography). The reconstruction time in PyTomography (5.4 s) was significantly faster than the time required in STIR (51542.2 s), owing to the speed-up factor of GPU implementation.
  2. 2.
    Clinical SPECT Validation With Vendor: A Jaszczak phantom was filled with Lu-177 and scanned at 180 projections ($128\times 128$ with a resolution of $4.8\text{mm}\times 4.8\text{mm}$) using a Siemens Symbia T Series SPECT/CT scanner with Medium Energy collimators. Reconstruction was performed with OSEM (4 iterations, 10 subsets) using (i) the manufacturer scanner software and (ii) PyTomography. Reconstructions and corresponding 1 dimensional profiles are shown in Figure 8; they appear similar. 3D slicer was used to obtain sample segmentations of the hot and cold regions; the mean source-to-background ratios were 20.97 (vendor) and 21.76 (PyTomography). The image noise, quantified by the standard deviation of predicted counts in the cold region, were 0.217 counts (vendor) and 0.178 counts (PyTomography).
  3. 3.
    Monte Carlo PET Validation With CASToR: The ultra-high resolution 18F-FDG PET/ T1-weighted MRI brain image34 used in the PET example (556 million events) was reconstructed using LM TOF OSEM (2 iterations, 21 subsets) with both PyTomography and CASToR. Resulting reconstructions are shown in Figure 9. The mean activity ratio in grey to white matter was 2.05 (PyTomography) and 2.03 (CASToR). The noise in the image was estimated as the standard deviation of counts in white matter divided by the mean activity in white matter; the values were 0.329 (PyTomography) and 0.321 (CASToR). Apart from processing the ROOT files and building the sensitivity image, the time required for reconstruction in PyTomography was 14 s to build the system matrix and required 59 s per iteration for reconstruction. CASToR reconstruction was ran on a separate HP Z8 Computer Desktop with 2 Intel Xeon Gold 6136 CPUs at 3.00 GHz with a total of 48 CPU cores. The reconstruction was performed with 24 CPU cores and reconstruction required approximately 455 s per full iteration.
The images shown in this section are displayed using the “nipy_spectral” colormap of matplotlib because it has significant contrast. Since this colormap is neither perceptually uniform nor monotonically increasing in lightness, images in this section should be viewed in color.
Figure 7: Central coronal slices (left) and 1D profiles (right) of reconstructed SIMIND SPECT projection data using OSEM (2it/8ss) with attenuation/PSF modeling; images were reconstructed using STIR and PyTomography. 1D profiles correspond to the horizontal lines shown on the coronal slices.
Figure 8: Sample axial slices (top) and 1D profiles (bottom) of a 177Lu Jaszczak phantom reconstructed using OSEM (4it/10ss) with attenuation/PSF modeling and triple energy scatter correction. Reconstructions were performed using (i) the scanner software, and (ii) PyTomography. 1D profiles correspond to the horizontal lines shown on the axial slices.
Figure 9: Sample coronal slices (top) and 1D profiles (bottom) of the high resolution brain phantom reconstructed using CASToR and PyTomography with LM TOF OSEM (2it/21ss). The two images were normalized such that the sum of their voxels was identical. 1D profiles correspond to the horizontal lines shown on the coronal slices
Acknowledgements
This work was supported by the Natural Sciences and Engineering Research Council of Canada (NSERC) CGS D Award 569711, NSERC Discovery Grants RGPIN-2019-06467 and RGPIN-2021-02965, as well as computational resources and services provided by Microsoft AI for Health.

### v4

@@
-PyTomography: A Python Library for Medical Image Reconstruction Luke Polson1,2, Roberto Fedrigo1,2, Chenguang Li1,2, Maziar Sabouri1,2, Obed Dzikunu2,3, Shadab Ahamed1,2, Nicolas Karakatsanis4, Arman Rahmim1,2,3,5,∗, Carlos Uribe2,5,6,∗
+PyTomography: A Python Library for Medical Image Reconstruction
 Abstract
-Background: There is an absence of open-source libraries in emission tomography that (i) use modern and popular backend code to encourage community contributions and (ii) offer support for the multitude of reconstruction algorithms available in recent literature, such as those that employ artificial intelligence.
-Purpose: To create and evaluate a GPU-accelerated, open-source, and user-friendly image reconstruction library, designed to serve as a central platform for the development, validation, and deployment of various tomographic reconstruction algorithms.
-Methods: PyTomography was developed using Python and inherits the GPU-accelerated functionality of PyTorch and parallelproj for fast computations.
-Its flexible and modular design decouples the system matrix from reconstruction algorithms, simplifying the process of integrating new imaging modalities using various python tools.
-Results: Example use cases demonstrate the software capabilities in parallel hole SPECT and listmode PET imaging.
-Use of the OSEM, BSREM, KEM, and DIPRecon reconstruction algorithms are explored.
-System matrices incorporating breathing-motion correction are explored in SPECT imaging.
-PyTomography is validated against alternative vendor/open source software for OSEM based reconstruction.
-Conclusions: We have developed and publicly shared PyTomography, a highly optimized and user-friendly software for medical image reconstruction, with a class hierarchy that fosters the development of novel imaging applications.
-I. Introduction
+There is a need for open-source libraries in emission tomography that (i) use modern and popular backend code to encourage community contributions and (ii) offer support for the multitude of reconstruction algorithms available in recent literature, such as those that employ artificial intelligence.
+The purpose of this research was to create and evaluate a GPU-accelerated, open-source, and user-friendly image reconstruction library, designed to serve as a central platform for the development, validation, and deployment of various tomographic reconstruction algorithms.
+PyTomography was developed using Python and inherits the GPU-accelerated functionality of PyTorch and parallelproj for fast computations.
+Its flexible and modular design decouples system matrices, likelihoods, and reconstruction algorithms, simplifying the process of integrating new imaging modalities using various python tools.
+Example use cases demonstrate the software capabilities in parallel hole SPECT and listmode PET imaging.
+Overall, we have developed and publicly share PyTomography, a highly optimized and user-friendly software for medical image reconstruction, with a class hierarchy that fosters the development of novel imaging applications.
+keywords:
+Medical Imaging , Image Reconstruction , SPECT , PET
+journal: SoftwareX
+Metadata
+Nr. Code metadata description Please fill in this column
+C1 Current code version v3.0.0
+C2 Permanent link to code/repository used for this code version https://github.com/qurit/PyTomography
+C3 Permanent link to Reproducible Capsule None. Reproducible examples provided as part of the documentation.
+C4 Legal Code License MIT License
+C5 Code versioning system used git
+C6 Software code languages, tools, and services used python
+C7 Compilation requirements, operating environments & dependencies python >=>=> = 3.11
+C8 If available Link to developer documentation/manual https://pytomography.readthedocs.io/en/latest/
+C9 Support email for questions https://pytomography.discourse.group/
+Table 1: Code metadata (mandatory)
+1 Motivation and significance
 Medical imaging forms a cornerstone in modern healthcare by providing visual and quantitative information about internal body structures and functions.
-It enables early detection1, accurate diagnosis, and precise treatment planning for a wide range of medical conditions2, 3, 4, contributing to improved patient outcomes and enhanced medical decision-making.
-Reconstruction algorithms are routinely used to generate 3D images that can be used for both research and clinical decision making 5.
-While the development and validation of reconstruction techniques remains an active field of research6, 7, 8, 9, it can often be difficult to share and disseminate one’s findings.
+It enables early detection [1], accurate diagnosis, and precise treatment planning for a wide range of medical conditions [2, 3, 4], contributing to improved patient outcomes and enhanced medical decision-making.
+Reconstruction algorithms are routinely used to generate 3D images that can be used for both research and clinical decision making [5].
+While the development and validation of reconstruction techniques remains an active field of research [6, 7, 8, 9], it can often be difficult to share and disseminate one’s findings.
 Additionally, while manufacturers of tomographic imaging equipment such as Single Photon Emission Computed Tomography (SPECT) and Positron Emission Tomography (PET) cameras provide their own internal reconstruction software, users cannot access all implementation details; this limits reproducibility in scientific studies.
@@
 A few tools have already been developed to try and address some of these issues.
-STIR 10 and CASToR 11 are two C/C++ based open source and collaborative image reconstruction frameworks that support a variety of reconstruction algorithms for SPECT and PET.
-NiftyPET 12 is a python library which employs GPU-accelerated code written in CUDA C for PET reconstruction.
-QSPECT13 and TIRIUS support use of the ordered subset expectation maximum (OSEM) algorithm for SPECT and PET reconstruction respectively.
-Other libraries, such as TomoPy 14 and the ASTRA toolbox 15, support a variety of GPU-accelerated image tomographic reconstruction algorithms, although they lack extensive SPECT/PET system modeling.
+STIR [10] and CASToR [11] are two C/C++ based open source and collaborative image reconstruction frameworks that support a variety of reconstruction algorithms for SPECT and PET.
+NiftyPET [12] is a python library which employs GPU-accelerated code written in CUDA C for PET reconstruction.
+QSPECT [13] and TIRIUS support use of the ordered subset expectation maximum (OSEM) algorithm for SPECT and PET reconstruction respectively.
+Other libraries, such as TomoPy [14] and the ASTRA toolbox [15], support a variety of GPU-accelerated image tomographic reconstruction algorithms, although they lack extensive SPECT/PET system modeling.
 Web links to all these tools, as well as a list of other available tools for image reconstruction, can be found on the NMMItools webpage of the Society of Nuclear Medicine and Molecular Imaging (SNMMI).
-The unique purpose of this research is to build upon recent efforts in GPU-accelerated reconstruction16, 17, 18, 19 and AI-based reconstruction7, 8, 9 to establish an open source medical image reconstruction programming framework written in a popular programming language that facilitates collaboration.
-Valuable past efforts have focused on CUDA programming to accelerate reconstruction times in SPECT16 and PET 16, 17 reconstruction through the use of multi-threading.
-However, potential contributors to an open-source framework may be unfamiliar with CUDA; any tool that prioritizes community contributions should employ familiar backend code.
-In Python, the library NumPy is widely known, but it lacks direct GPU acceleration support.
-PyTorch, however, emerges as a suitable alternative since it maintains a similar interface to NumPy while also facilitating GPU acceleration and direct interface with neural networks.
-Some potential contributors may wish to provide contributions using GPU-accelerated functionality offered by other libraries, such as CuPy; an optimal framework should thus also permit the ability to incorporate alternative implementations for particular tasks.
-To this end, the python library PyTomography has been developed with the following priorities: (i) implementation of standard and traditional imaging modalities and reconstruction algorithms, (ii) implementation and dissemination of recent research developments, such as the deep image prior8, and (iii) encouragement of community involvement via extensive documentation and user-friendly tutorials.
-While the present focus of PyTomography is the development and validation of reconstruction algorithms for SPECT/PET, implementation of other imaging modalities and reconstruction algorithms can be added using the building blocks provided.
+The unique purpose of our present research and effort is to build upon recent efforts in GPU-accelerated reconstruction [16, 17, 18, 19] and AI-based reconstruction [7, 8, 9] to establish an open source medical image reconstruction programming framework written in a popular programming language that facilitates collaboration.
+Valuable past efforts have focused on CUDA programming in languages such as C/C++ to accelerate reconstruction times in SPECT [16] and PET [16, 17] reconstruction through the use of multi-threading.
+While some contributors to an open-source framework may prefer to use C/C++ programming, others may be unfamiliar with these languages; any tool that prioritizes community contributions should permit both options.
+Some of the advantages of a python framework are that (i) it is a widely known programming language and thus encourages community involvement and (ii) it permits the embedding of C/C++ code for particular applications.
+Python can directly utilize CUDA acceleration via the PyTorch library, for example, allowing users unfamiliar with C/C++ code to achieve GPU acceleration.
+To this end, we developed the python library PyTomography with the priorities to: (i) implement standard and traditional imaging modalities and reconstruction algorithms, (ii) implement and disseminate of recent research developments, such as the deep image prior [8], and (iii) encourage community involvement via extensive documentation and user-friendly tutorials.
+Based on the motivations previously described, the backend of library has been developed using PyTorch because it maintains a similar functional structure to the widely used NumPy library.
+Other libraries are used for more particular tasks; the python library paralleproj [19], which was developed with CUDA C backend code, is used for line integral computation in PET projectors.
+In general, code written in any language can be integrated in PyTomography provided a python interface is available.
+While the present focus of PyTomography is the development and validation of reconstruction algorithms for SPECT and PET, implementation of other imaging modalities and reconstruction algorithms can be added using the building blocks provided.
 Overall, the aim of this study is to introduce the community to the PyTomography package, demonstrate its current capabilities and flexibility, and to encourage community involvement in the continuation of its development.
-To demonstrate the versatility of the software, example applications will be explored in (i) SPECT reconstruction using PyTorch based rotate+sum projectors, and (ii) listmode (LM) time of flight (TOF) PET reconstruction using Joseph’s projectors from the python library parallelproj19.
-In what follows, we elaborate on our methods and results for PyTomography, which is made publicly available to download; relevant links can be found in section VI..
-II. Materials and Methods
-II.A. Mathematical Description
+To demonstrate the versatility of the software, example applications are explored in (i) SPECT reconstruction using PyTorch based rotate+sum projectors, and (ii) listmode (LM) time-of-flight (TOF) PET reconstruction.
+In what follows, we elaborate on our methods and results for PyTomography, which is made publicly available to download.
+2 Software description
+2.1 Software scope and functionality
 A brief overview of the mathematical paradigm of medical imaging is presented in order to motivate the software architecture.
@@
 Components of a vector are given by $v_{i}$, and components of linear operators are given by $A_{ij}$.
-II.A.1. Projectors
-In the paradigm of tomographic medical imaging, there are two main vector spaces: the “projection” space $\mathcal{V}$ consisting of measured projection data $g$ and the “object” space $\mathcal{U}$ consisting of 3D objects $f$.
+When $A1$ or $A^{T}1$ is written, “1” is to be interpreted as a vector (in the appropriate vector space) with all elements equal to 1.
+2.1.1 Projectors
+In the paradigm of tomographic medical imaging, there are two main vector spaces: the “projection” space $\mathcal{V}$ consisting of measured data $g$ and the “object” space $\mathcal{U}$ consisting of 3D objects $f$.
 Since measured and reconstructed data is digitized, $\mathcal{U}$ and $\mathcal{V}$ are assumed here to be finite dimensional vector spaces.
 In SPECT imaging, for example, $\mathcal{V}$ consists of a set of 2D pixelated images acquired by counting photons emitted by radiopharmaceuticals, and $\mathcal{U}$ represents the voxelized spatial distribution of radioactivity concentration in the object being imaged.
-Certain imaging systems can be characterized by a system matrix $H$, which corresponds to a linear operator mapping from object space to projection space.
-Its components $H_{ij}$ represent a contribution of voxel $j$ to detector element $i$.
-In SPECT/PET, analytical modeling of the data expectation $\bar{g}$ can be achieved via $\bar{g}=Hf+s$ where $s$ represents scatters and randoms.
+Linear imaging systems can be characterized by a system matrix $H:\mathcal{U}\to\mathcal{V}$.
+Its components $H_{ij}$ represent the contribution of element $j$ in object space to element $i$ in projection space.
+In emission tomography, analytical modeling of the data expectation $\bar{g}$ can be achieved via $\bar{g}=Hf+s$ where $s$ is an additive term containing estimates for scattered events (SPECT) or scatter and random coincidences (PET).
 In CT, analytical modeling is achieved via $\bar{g}=be^{-Hf}$ where $b$ denotes the rate of emissions from the transmission source to a particular measurement.
 The operation of $H$ on an object vector is known as forward projection (FP), while the operation of $H^{T}$ on a projection vector is known as back projection (BP).
-Efficient implementation of FP and BP depends on the particular imaging modality; this typically involves decomposing $H$ into a sequence of less computationally burdensome operations.
-In general, FP can be decomposed into the following sequence: (i) linear endomorphisms mapping from $\mathcal{U}\to\mathcal{U}$, (ii) operators mappings from $\mathcal{U}\to\mathcal{V}$, and (iii) linear endomorphisms mapping from $\mathcal{V}\to\mathcal{V}$ given by $\{B_{k}\}$.
-BP is given by the transpose of this operation sequence.
-In PyTomography, operators which map to the same vector space are defined as transforms, while operators that map between different vector spaces are defined as projectors.
 Example implementations available in PyTomography are listed below:
-  • Parallel Collimator SPECT: Modeling of primary detections in this imaging modality can be approximated using the “rotate+sum” technique16, whereby (i) consists of rotations and attenuation/collimator detector response modeling, (ii) consists of summation along an axis, and (iii) consists of adjustments for a camera with a finite field of view.
-System modeling is approximated using the following system matrix:
+  • 1.
+    Parallel Collimator SPECT: Modeling of primary detections in this imaging modality can be approximated using the “rotate+sum” technique [16]; in this case, the system matrix can be written as
     $H=\sum_{n}v_{n}\otimes\left(\mathcal{T}_{\text{fov}}^{(n)}~{}\mathcal{P}_{x}~{}\mathcal{T}_{\text{cdr}}^{(n)}\mathcal{T}_{\text{att}}^{(n)}~{}\mathcal{T}_{\text{trans}}^{(n)}~{}\mathcal{T}_{\text{rot}}^{(n)}\right)$ (1)
-    where $n$ designates the projection angle, $\mathcal{T}_{\text{rot}}^{(n)}$ (transform) implements a rotation, $\mathcal{T}_{\text{trans}}^{(n)}$ (transform) implements a translation that accounts for non-central projections, $\mathcal{T}_{\text{att}}^{(n)}$ (transform) implements the adjustment for photon attenuation, $\mathcal{T}_{\text{cdr}}^{(n)}$ (transform) implements depth-dependent collimator blurring, $\mathcal{P}_{x}$ (projector) is summation along the $+x$ axis, and $\mathcal{T}_{\text{fov}}^{(n)}$ (transform) applies a mask in projection space that accounts for a finite field of view, $v_{n}$ is a unit vector of length $N$ with 1 in component $n$, and $\otimes$ is the outer product.
-  • LM PET: Modeling in this modality can be approximated when (i) consists of object-space resolution modeling, (ii) consists of ray tracing between all detected lines of response (LOR), and (iii) consists of normalization/attenuation correction for each LOR.
-In this case, the system matrix can be written as
+    where $n$ designates the projection angle, $\mathcal{T}_{\text{rot}}^{(n)}$ implements a rotation, $\mathcal{T}_{\text{trans}}^{(n)}$ implements a translation that accounts for non-central projections, $\mathcal{T}_{\text{att}}^{(n)}$ implements the adjustment for photon attenuation, $\mathcal{T}_{\text{cdr}}^{(n)}$ implements depth-dependent collimator blurring, $\mathcal{P}_{x}$ is summation along the $+x$ axis, and $\mathcal{T}_{\text{fov}}^{(n)}$ applies a mask in projection space that accounts for a finite field of view, $v_{n}$ is a unit vector of length $N$ with 1 in component $n$, and $\otimes$ is the outer product.
+  • 2.
+    PET: Modeling of primary events (no randoms or scatters) in PET can be approximated as
     $H=\mathcal{T}_{\eta}~{}\mathcal{T}_{\omega}~{}\mathcal{P}_{\Sigma}~{}\mathcal{T}_{rr}$ (2)
-    where $\mathcal{T}_{rr}$ (transform) is used for object-space resolution modeling, $\mathcal{P}_{\Sigma}$ (projector) is used to compute line integrals, $\mathcal{T}_{\eta}$ (transform) adjusts for detector normalization, and $\mathcal{T}_{\omega}$ (transform) adjusts for attenuation effects.
-Due to cancellation in certain reconstruction algorithms, $\mathcal{T}_{\eta}$ and $\mathcal{T}_{\omega}$ need not always be included (though they are required for computing the sensitivity factor).
+    where $\mathcal{T}_{rr}$ implements object-space resolution modeling (typically via a uniform Gaussian kernel), $\mathcal{P}_{\Sigma}$ implements line integral computation (listmode/sinogram), $\mathcal{T}_{\eta}$ adjusts for detector normalization, and $\mathcal{T}_{\omega}$ adjusts for attenuation effects.
 PyTomography can be used to construct projectors that extend projection space; such functionality is necessary, for example, when modeling multiple photopeaks in SPECT or when using motion correction.
@@
 $H^{\prime}=\sum_{n}v_{n}\otimes B_{n}H_{n}A_{n}$ (3)
-where $\left\{A_{n}\right\}$ are object-to-object space transforms, $\left\{H_{n}\right\}$ are a sequence of system matrices, $\left\{B_{n}\right\}$ are a sequence of projection-to-projection space transforms, $v_{n}$ is a basis vector of length $N$ with a value of $1$ in component $n$ and $\otimes$ is an outer product.
+where $\left\{A_{n}\right\}$ and $\left\{B_{n}\right\}$ are additional operations, $\left\{H_{n}\right\}$ are a sequence of system matrices, $v_{n}$ is a basis vector of length $N$ with a value of $1$ in component $n$, and $\otimes$ is an outer product.
 This can be utilized in the following examples:
-  • Periodic Motion Correction: $A_{n}$ represent spatial transformation operators corresponding to time gate $n$, $H_{n}$ represent system matrices for each motion transform (attenuation modeling may depend on object deformation), and $B_{n}=1$.
+  • 1.
+    Periodic Motion Correction: $A_{n}$ represent spatial deformation operators corresponding to time gate $n$, $H_{n}$ represent system matrices for each motion transform (attenuation modeling may depend on object deformation), and $B_{n}=1$.
 This maps to an extended projection space $(N,...)$ where $N$ is the number of different time gates (e.g.
 during a breathing cycle).
+  • 2.
     Multiple SPECT Photopeaks: In this scenario, $A_{n}$ are scaling factors that adjust for detector sensitivities and branching ratios, $B_{n}$ are identity mappings, and $H_{n}$ represents the system matrix corresponding to a particular photopeak.
 The extended projection space has dimensions $(N,...)$ where $N$ is the number of different photopeaks.
-II.A.2. Likelihoods
-A likelihood function $L(g|\bar{g})$ provides the likelihood of observing data $g$ given expectation $\bar{g}$.
-Since $\bar{g}$ is approximated via analytical models depending on $f$ and $H$, it follows that $L(g|\bar{g})=L(g|f,H,...)$, where the ellipsis “...” contains additional parameters such as a scatter estimate $s$; the ellipsis will assumed implicitly for the remainder of this paper.
-Many iterative reconstruction algorithms seek to maximize the likelihood function via gradient ascent using the gradient $\nabla_{f}L(g|f,H)$.
+2.1.2 Likelihoods
+A statistical likelihood function $\tilde{L}(g|\bar{g})$ is proportional to the probability of observing an actual realization of data $g$ given expectation $\bar{g}$.
+Since $\bar{g}$ is approximated using $f$ and $H$, it follows that $\tilde{L}(g|\bar{g})=\tilde{L}(g|f,H,...)$, where the ellipsis “...” contains additional parameters such as a scatter estimate $s$; the ellipsis will be assumed implicit for the remainder of this paper.
+Many iterative reconstruction algorithms seek to maximize the likelihood function via gradient ascent by exploiting the gradient $\nabla_{f}\tilde{L}(g|f,H)$.
+Since (i) computation of $\nabla_{f}\log\tilde{L}(g|f,H)$ is more tenable than $\nabla_{f}\tilde{L}(g|f,H)$ and (ii) maximization of $\log\tilde{L}(g|f,H)$ is equivalent to maximization of $\tilde{L}(g|f,H)$, the statistical “log-likelihood” is nearly always considered instead.
+In PyTomography, the “likelihood” $L(g|f,H)$ refers to an objective function to be maximized; as such, the “likelihood” $L(g|f,H)$ is often used interchangeably with the statistical “log-likelihood” $\log\tilde{L}(g|f,H)$ for a particular modality.
 Likelihood functions for different modalities are shown in Table 2.
 Table 2: Likelihoods and corresponding gradients for various medical imaging modalities; $s$ represents a scatter estimate used in SPECT/PET while $b$ represents a scaling factor used in CT.
+The CT likelihood corresponds to an ideal scenario where the number of counts in each sinogram bin follows a Poisson distribution; this is not necessarily the case in real systems with complex electronic noise.
 Likelihood $L(g|f,H)$ $\nabla_{f}L(g|f,H)$
@@
 Weighted MSE $-\frac{1}{2}\sum_{i}\left[W\left(g-Hf\right)^{2}\right]_{i}$ $H^{T}W(g-Hf)$
-II.A.3. Reconstruction Algorithms
+2.1.3 Reconstruction Algorithms
 An image reconstruction algorithm $A$ uses measured data $g$ (i.e.
-projections) to estimate a corresponding object $f$ that would produce projections $g$ given a system matrix $H$.
-Statistical algorithms find the image estimate $f$ which maximizes a likelihood function $L(g|H,f)$ characterizing the imaging system.
+projections) to estimate a corresponding object $f$ that would produce this $g$ given a system matrix $H$.
+Statistical algorithms find the image estimate $f$ which maximizes the likelihood function $L(g|H,f)$ characterizing the imaging system.
 This can be expressed as
@@
 The main algorithms in PyTomography are preconditioned gradient ascent (PGA) algorithms; these algorithms aim to maximimize the likelihood via iterative updates that take the form
-$f^{n+1}=f^{n}+C^{n}(f^{n},H_{n})\left[\nabla_{f}L(g^{n}|f^{n},H_{n})-\beta_{n}\nabla_{f}V(f^{n})\right]$ (5)
-where $n$ is the subiteration index, $f^{n}$ is the object estimate of subiteration $n$, $g^{n}$ specifies a subset of data used in subiteration $n$, $H_{n}$ is a system matrix projecting only to the current subset, $C^{n}(f^{n})$ is a preconditioner factor that depends on the object estimate, $V(f^{n})$ is a prior function used to regularize the algorithm, and $\beta_{n}=c_{n}\beta$ is a scaling factor for the prior function (scaled by the fraction of subset data $c_{n}$).
+$f^{n+1}=f^{n}+C^{n}(f^{n},H^{n})\left[\nabla_{f}L(g^{n}|f^{n},H^{n})-\beta^{n}\nabla_{f}V(f^{n})\right]$ (5)
+where $n$ is the subiteration index (used as a superscript and not an exponent for all variables), $f^{n}$ is the object estimate of subiteration $n$, $g^{n}$ specifies a subset of data used in subiteration $n$, $H^{n}$ is a system matrix projecting only to the current subset, $C^{n}(f^{n})$ is a preconditioner factor that depends on the object estimate, $V(f^{n})$ is a prior function used to regularize the algorithm, and $\beta^{n}=c^{n}\beta$ is a scaling factor for the prior function (scaled by the fraction of subset data $c^{n}$).
 These algorithms partition the data $g$ into $M$ approximately equal subsets; a single iteration of an algorithm is defined as the number of subiterations where all subsets have been used once.
 Various PGA algorithms available in PyTomography are shown in Table 3; they are characterized by their preconditioner factor.
-Additional reconstruction techniques can be developed using these algorithms.
-The kernel expectation maximum (KEM), for example, corresponds to (i) reconstructing a coefficient image $\alpha$ using $\alpha=A_{\text{OSEM}}(g,HK)$ where $K$ is a kernel transform and (ii) obtaining the image estimate via $f=K\alpha$.
-Table 2: Sample algorithms available in PyTomography: shown are ordered subset expectation maximum (OSEM), ordered subset maximum a posteriori one step late (OSMAPOSL), block sequential regualized expectation maximum (BSREM), and kernelized expectation maximum (KEM).
-Algorithm Preconditioner factor $C^{n}(f^{n},H)$
-OSEM20 $f^{n}/\left(H_{n}^{T}1\right)$ and $V(f^{n})=0$
-OSMAPOSL21 $f^{n}/\left(H_{n}^{T}1+\nabla_{f}V(f^{n})\right)$
-BSREM22 $f^{n}\alpha(n)/\left(\omega_{n}H^{T}1\right)$
 Certain algorithm names in the literature correspond to specific realizations of the generic algorithms in Table 3 with particular likelihoods in Table 2.
 The OS-SART algorithm, for example, corresponds to the use of OSEM with the weighted MSE likelihood (with weights $W=1/(H1)$).
+Table 3: Sample algorithms available in PyTomography: shown are ordered subset expectation maximum (OSEM), ordered subset maximum a posteriori one step late (OSMAPOSL), block sequential regualized expectation maximum (BSREM), and kernelized expectation maximum (KEM).
+$n$ is used as a superscript (not an exponent) for all variables.
+For BSREM, $\alpha^{n}$ is a relaxation parameter.
+Algorithm Preconditioner factor $C^{n}(f^{n},H^{n})$
+OSEM [20] $f^{n}/\left(\left(H^{n}\right)^{T}1\right)$ and $V(f^{n})=0$
+OSMAPOSL [21] $f^{n}/\left(\left(H^{n}\right)^{T}1+\nabla_{f}V(f^{n})\right)$
+BSREM [22] $f^{n}\alpha^{n}/\left(c^{n}H^{T}1\right)$
 Prior functions $V(f)$ used in reconstruction often take the form
@@
 When $w_{rs}$ is non-zero only for neighbouring voxels, the prior is known as a nearest neighbour prior; various nearest neighbour priors implemented in PyTomography are shown in Table 4.
+While the weighting factor $w_{rs}$ is typically proportional to inverse Euclidean distance, it is also possible to allow it to depend on an additional anatomical image $\chi$ whereby $w_{rs}=w_{rs}(\chi_{r},\chi_{s})$ Since anatomical information is used, this is commonly referred to as an anatomical prior (AP); variations of this technique have recently become popular in clinical practice due to enhanced lesion quantitation and detectability in bone SPECT/CT [6].
+In this paper, use of an AP refers to using only the 8 most similar neighbours based on absolute differences of HU values in a CT scan or attenuation map.
 Table 4: Sample nearest neighbour functions available in PyTomography; these priors are characterized by their image weighting function $\phi(f_{r},f_{s})$.
@@
 Log-cosh $\tanh((f_{r}-f_{s})/\delta)$
-Relative difference penalty (RDP) 24 $\left((f_{r}-f_{s})^{2}\right)/\left(f_{r}+f_{s}+\gamma|f_{r}-f_{s}|\right)$
-While the weighting factor $w_{rs}$ is typically proportional to inverse Euclidean distance, it is also possible to allow it to depend on an additional anatomical image $\chi$ whereby $w_{rs}=w_{rs}(\chi_{r},\chi_{s})$ Since anatomical information is used, this is commonly referred to as an anatomical prior (AP); variations of this technique have recently become popular in clinical practice due to enhanced lesion quantitation and detectability in bone SPECT/CT6.
-In this paper, use of an AP refers to using only the 8 most similar neighbours based on absolute differences of HU values in a CT scan or attenuation map.
+Relative difference penalty (RDP) [23] $\left((f_{r}-f_{s})^{2}\right)/\left(f_{r}+f_{s}+\gamma|f_{r}-f_{s}|\right)$
 PyTomography also offers support for the Deep Image Prior Reconstruction (DIPRecon) algorithm.
 Details regarding the implementation of this algorithm can be found in Algorithm 2 of Gong et.
-al8.
+al [8].
 The purpose of this reconstruction technique is to use anatomical images (CT/MR) to enhance functional images (SPECT/PET) through a sequence of neural network overfitting and mixing with OSEM updates.
@@
 This paper explores the use of DIPRecon on an ultra-high resolution PET/MR brain phantom.
-II.B. Software Architecture
-The key points of the last section are that (i) medical imaging consists of many different modalities that each require unique implementations for system modeling and (ii) tomographic reconstruction algorithms are modality agnostic.
-PyTomography has been carefully designed to permit arbitrary implementations of system modeling that can interface with the available reconstruction algorithms.
-The components of PyTomography, designed to facilitate the architecture described, are listed below.
+2.2 Software architecture
+PyTomography has been carefully designed to permit arbitrary implementations of system matrices (projectors) which can be used in various likelihood functions that themselves interface with reconstruction algorithms.
+The components of PyTomography were designed to facilitate this architecture and are listed below.
+A visual representation of the software library along with sample code representing a SPECT reconstruction use case is shown in Figure 1.
   1. 1.
-    Data. The standard data type used to represent images and raw projection data is the tensor class of PyTorch.
+    Data/Metadata.
+The standard data type used to represent images and raw projection data is the tensor class of PyTorch; data can be stored on CPU or GPU.
+Metadata classes contain all auxiliary information needed for image reconstruction pertaining to a particular imaging system, such as voxel size, spatial dimensions, and information pertaining to the raw data of a particular modality.
   2. 2.
-    Metadata.
-Metadata classes contain all auxiliary information needed for image reconstruction pertaining to a particular imaging system, such as voxel size, spatial dimensions, and information pertaining to the raw data of a particular modality.
+    Projectors.
+Projectors are mathematical operators that map between object and projection space.
+The system matrix is an example of a projector that models a particular imaging system.
+While the input/output datatype to projectors is the PyTorch tensor class, internal code can utilize functionality from other libraries.
+Projector subclasses inherit from a generic “SystemMatrix“ class and implement a “forward”, “backward”, and “compute_normalization_factor” method.
+Projectors used in subset based algorithms should additionally implement “set_n_subsets”, “get_projection_subset” and ”get_weighting_subset”; the second method takes in projection data and returns subset data based on how the projector defines subsets, and the third gets the corresponding weight ($c^{n}$) for scaling preconditioners.
   3. 3.
-    Transforms.
-Transform classes are used as utilities to model various features of the imaging process, such as attenuation and image blurring in SPECT.
-They are commonly used as components in a system matrix.
-Specifically, transforms are defined by the mathematical property that they are linear endomorphisms.
-  4. 4.
-    Projectors.
-Unlike transforms, projectors are mathematical operators that map between different vector spaces.
-The system matrix, which maps from object space to projection space and represents a model of the imaging system, is an example of a projector.
-While the input/output datatype to projectors is the PyTorch tensor class, internal code can utilize functionality from other libraries.
-  5. 5.
     Likelihoods.
@@
 Likelihoods $L(g|H,f)$ are initialized using a system matrix $H$, projection data $g$, and any additional dependencies; once initialized they act as functions of an image estimate $f$.
-  6. 6.
+Likelihoods inherit from a generic “Likelihood” class and should implement the “compute_gradient” method.
+Higher order derivatives of the likelihood may also be defined as seperate methods within the class.
+  4. 4.
     Algorithms.
@@
 Statistical iterative algorithms require as input a likelihood instance, and may take as optional arguments prior functions and callbacks.
-A visual representation of the software library along with sample code representing a SPECT reconstruction use case is shown in Figure 1.
-Metadata is first obtained via the raw vendor data stored in the path represented by the variable “file_NM”.
-The metadata is used as input to a “SPECTSystemMatrix” class, which also takes as input attenuation, PSF, and field of view cutoff transforms (creation of transforms not shown).
-The generated system matrix is used to build a likelihood instance, which also takes in as input the photopeak (creation not shown) and scatter as an additive term (creation not shown).
-The likelihood is then used to initialize an OSEM class for reconstruction; a reconstructed object is subsequently obtained using 4 iterations and 8 subset.
-Figure 1: Basic overview of the workflow of PyTomography (top) with code representing a specific SPECT reconstruction use case (bottom).
-Sample classes and implementations are shown; these do not comprise the full set of classes available in the software.
-II.C. Examples
+Unlike projectors and likelihoods, algorithms do not inherit from a generic base class.
+In particular, however, linear PGAAs inherit from a generic “LinearPreconditionedGradientAscentAlgorithm” class and simply need to implement the “_linear_preconditioner_factor” and optionally reimplement the “__init__” method to take in required arguments.
+PyTomography also has an input/output (I/O) module for reading various file formats; it currently supports SIMIND, DICOM (SPECT/CT), GATE (PET), and HDF5 listmode (PET) formats.
+Data reading consists of loading the generated files and constructing the appropriate data/metadata instances within the PyTomography framework.
+The I/O module is also used to read internal files of the library, such as coefficients to convert Hounsfield Units to attenuation coefficients, files containing SPECT collimator information, and files containing PET scanner geometry.
+Novel development of a projector with associated data sources should thus consist of the following (i) implementation of metadata classes, (ii) I/O code to load data/metadata from the data sources, and (iii) development of a projector to interface with the data/metadata.
+Metadata class instances may be initialized by various data sources; for example, the same “SPECTProjMeta” class is initialized from SIMIND or DICOM I/O data loading functions, even though they represent different data sources.
+Metadata is thus linked to a modality and not to a specific data source.
+Deployment of the software was designed to maximize accessibility.
+In addition to placing the backend code on GitHub, the library was also uploaded to the Python Package Index (PyPi) so users can easily install the library using the simple “pip install pytomography” terminal command.
+A readthedocs documentation website was created that also contains a comprehensive set of tutorials in jupyter notebook format.
+While many of the tutorials are aimed at training end users, there also exist developer tutorials for implementing new projectors in the software.
+A discourse website was developed for two purposes: (i) to allow end users to ask questions about how to use the software and (ii) to permanently store and make public the questions/answers as they may be useful for other users of the library.
+Figure 1: Basic overview of the workflow of PyTomography when using statistical algorithms (left) with code representing a specific SPECT reconstruction use case (right).
+Shown are the minimum required inputs to the “__init__” function.
+the minimum required methods for each class, and sample subclasses available in the library.
+The “…” represents additional arguments that may be present in subclasses.
+In the SPECT use case, for example, the additional input argument “obj2obj_transforms” is implemented to take in operations that implement attenuation and PSF modeling.
+Not shown in the sample code: importing required libraries, loading the photopeak/scatter data, creating the attenuation/PSF transforms, and specifying the filepath “file_NM”.
+3 Illustrative examples
 Multiple examples are explored to highlight the capabilities of PyTomography.
-At present, reconstruction of parallel collimator SPECT and LM PET data is demonstrated, but it should be emphasized that these are purely intended as example use cases for particular imaging modalities.
+Specifically, reconstruction of parallel-collimator SPECT and LM-TOF PET data is demonstrated, but it should be emphasized that these are only intended as example use cases for particular imaging modalities.
+Validation against other software libraries is also shown in the Appendix, which contains (i) SIMIND simulated SPECT data reconstruction validated against STIR, (ii) DICOM SPECT data (obtained on a GE Tandem Discovery 670 Pro) validated against MIM’s SPECT reconstruction functionality, and (iii) GATE PET data reconstruction validated against CASToR.
+We chose to validate DICOM reconstruction against MIM, and not the Siemens internal reconstruction software, because it represents an external library with access to the same information as PyTomography.
 All computations were performed using a Microsoft Azure virtual machine (Standard NC6s v3) with a 6 CPUs (Intel(R) Xeon(R) CPU E5-2690 v4 @ 2.60GHz), 112 GB of RAM, and a TeslaV100 GPU.
 Python scripts and notebooks used to obtain these results can be found at https://github.com/qurit/PyTomography_paper_code.
-II.C.1. Monte Carlo SPECT
-Two use cases are explored for reconstruction of Monte Carlo generated SPECT image data from the SIMIND Monte Carlo simulation program26.
-The system matrices used in each case correspond to equation 1, with each employing attenuation modeling and PSF modeling.
-The triple energy window technique27 is used for scatter correction.
-[…]
-Table 4: Reconstruction parameters used for SPECT reconstruction in this paper.
-Neighbourhood size corresponds to the kernel size in KEM, and prior neighbourhood for BSREM.
-Nearest neighbours were obtained from attenuation maps.
-Algorithm OSEM BSREM KEM
-Subsets 8 8 8
-Neighbourhood Size - $3\times 3\times 3$ $5\times 5\times 5$
-$k$ Nearest Neighbours - 8 40
-Kernel - - Eq. 5 of Vuohijoki et. al. 6
-Prior - RDP-AP -
-[…]
-II.C.2. Clinical SPECT
-This example highlights SPECT reconstruction on publicly available data from the Deep Blue data repository31 that was used for the SNMMI 177Lu dosimetry challenge32, 33.
-Patients received radiopharmaceutical therapy with 177Lu-DOTATATE for neuroendocrine tumours (4 cycles of 7.4 GBq/cycle administered every 8 weeks); scans were taken at 4 different time points in the week following a therapeutic dose.
-Images consisted of 120 projections of shape $256\times 256$ with resolution $2.4~{}\text{mm}\times 2.4~{}\text{mm}$.
-The system matrix employed attenuation, PSF, and finite FOV modeling.
-Reconstructions using OSEM, BSREM, and KEM were performed for 50 iterations each; the TEW technique was used to correct for scatter in all cases.
-II.C.3. Monte Carlo PET
-This example demonstrates (i) the capabilities of LM TOF PET reconstruction on data generated by the GATE Monte Carlo simulation program and (ii) the DIP Recon technique.
-The simulated scanner consisted of $8$ rings of 338 mm radius with 32.25 mm ring spacing, and a TOF coincidence resolution of 550 ps.
-For simplicity, data acquisition in the simulation ignored dead-time effects and only considered true primary coincidence events; coincidences due to phantom-scatter and randoms were ignored.
-The digital phantom used in simulation corresponded to an ultra-high resolution 18F-FDG PET/ T1-weighted MRI brain image34; the full simulated dataset was representative of a 3 min scan and consisted of $556$ million coincidence events.
-Data was reconstructed using a $128\times 128\times 96$ matrix size with voxel spacing of 222~{}2mm.
+3.1 SPECT
+This section highlights SPECT reconstructions using available reconstruction algorithms in PyTomography, demonstrating differences in quantitative metrics between different algorithms.
+Two examples are considered: (i) reconstruction of Monte Carlo SPECT data generated via SIMIND, and (ii) reconstruction of clinical SPECT data.
+In the clinical case, the patient received radiopharmaceutical therapy with 177Lu-PSMA for prostate cancer and was imaged 3 hours post injection; the XCAT phantom used as input for SIMIND was configured such that the kidney, liver, and lesion activity concentrations roughly approximated the clinical patient.
+In each case, the data correspond to acquisition on a GE Tandem Discovery 670 Pro SPECT system (MEGP collimators) with 60 projections (25 s per projection).
+The clinical data consists of acquisition at two bed positions with projections of shape $128\times 128$, while the simulated data consists of acquisition at one bed position with projections of shape $128\times 240$.
+The two lesions inserted in the XCAT phantom were (i) based on an observed bone metastasis in the sternum of the clinical patient, and (ii) placed in the liver.
+In each case, the system matrix employed attenuation and PSF modeling, and the likelihood function considered an additive scatter term obtained via the triple energy window technique.
+The reconstruction algorithms considered were OSEM (4it/6ss), BSREM-RDP (40it/6ss), BSREM-RDPAP (40it/6ss), and KEM (10it/6ss) [’it’ for iterations; and ’ss’ for subsets]; algorithms using RDP used the hyperparameters $\beta=0.3$, $\gamma=2$, RDPAP only considered the 8 nearest neighbours based on the attenuation map, and KEM used Eq.
+5 of Vuohijoki [6] for its filter.
+The time required for reconstruction of one bed position of the clinical data was 1.07 s/it (OSEM), 0.83 s/it (BSREM-RDP), 0.86 s/it (BSREM-RDPAP), and 1.07 s/it (KEM).
+The time required for reconstruction of the simulated data was 1.32 s/it (OSEM), 1.13 s/it (BSREM-RDP), 1.21 s/it (BSREM-RDPAP), and 1.52 s/it (KEM).
+For the clinical data, the left kidney was segmented using the CT-based TotalSegmentator model [24], and the bone metastasis in the sternum was segmented by first segmenting the lesion from the OSEM reconstruction, and then applying a threshhold mask greater than 0.150.150.15~{}0.15cm-1 so that only voxels within the bone were considered.
+A spherical mask was used in the liver for both the simulated and clinical data for measurement of noise in each reconstruction algorithm; noise was quantified as a percent by computing the standard deviation of uptake divided by the mean of uptake in that region.
+Reconstructions of the clinical and simulated data are shown in Figures 2 and 3 respectively.
+Figure 2: Top: Maximum intensity projections of reconstructed clinical SPECT data; the leftmost subplot corresponds to reconstruction using the MIM reconstruction software library.
+Three annotations are shown: (i) the left kidney, contained inside a pink box, with the corresponding activity concentration in kBq/mL (ii) the location of the spherical mask in the liver used for noise quantification with the corresponding noise in percent, and (iii) the location of the bone metastasis (blue arrow) with the corresponding activity concentration in kBq/mL.
+Bottom: Axial slice corresponding the hottest slice of the bone metastasis; SPECT uptake is shown in color and the attenuation map is shown in greyscale.
+Figure 3: Top: Maximum intensity projections of reconstructed simulated SPECT data; the leftmost subplot corresponds to the ground truth XCAT phantom used for simulation in SIMIND.
+Four annotations are shown: (i) the left kidney, contained inside a pink box, with the corresponding activity concentration in kBq/mL (ii) the location of the spherical mask in the liver used for noise quantification with the corresponding noise in percent, and (iii) the location of the bone metastasis (blue arrow) with the corresponding activity concentration in kBq/mL, and (iv) the location of the liver metastasis (purple arrow) with the corresponding activity concentration in kBq/mL.
+Bottom: Axial slice corresponding the hottest slice of the bone metastasis; SPECT uptake is shown in color and the attenuation map is shown in greyscale.
+3.2 PET
+PyTomography is capable of reconstructing real and simulated PET data in either sinogram mode or listmode.
+This example uses GATE Monte Carlo data and demonstrates the following capabilities of the library in PET reconstruction:
+  1. 1.
+    LM-TOF reconstruction
+  2. 2.
+    Random estimation via delayed coincidences and scatter estimation using the TOF single scatter simulation (TOF-SSS) technique
+  3. 3.
+    Use of the DIPRecon algorithm to enhance reconstruction
+The simulated scanner consisted of $8$ rings of 338 mm radius with 32.25 mm ring spacing, and a TOF coincidence resolution of 550 ps; this fictional scanner is representative of a Siemens Biograph mMR with TOF capabilities.
+The digital phantom used for simulation in GATE consisted of an ultra-high resolution 18F-FDG PET image, a T1-weighted MRI brain image, and a corresponding attenuation map [25].
+The total uptake of 15.66 MBq was representative of a typical 18F-FDG distribution 1 hour post injection.
 The normalization $\eta$ was obtained via simulation of a thin cylindrical shell of radius 318318318~{}318mm that extended along the axial FOV of the scanner; normalization computation is available in PyTomography for basic polygonal PET geometries.
-Reconstruction was performed with normalization correction in 3 different ways:
+In all cases, simulated data were reconstructed using a $128\times 128\times 96$ matrix size with voxel spacing of 222~{}2mm; 5 TOF bins were used with a total range of 30 cm.
+Before reconstruction, the “randoms” contribution to each listmode event was estimated from delayed coincidences; this process consisted of (i) binning the delayed coincidence events into a sinogram (ii) smoothing the sinogram to reduce noise, and (iii) converting back into LM form for use in LM reconstruction.
+The “scatters” contribution for each listmode event was estimated using the TOF single scatter simulation (SSS) capabilities of the library, based on Watson [26].
+The PET image used as a proxy for SSS was a reconstruction of the data via LM-TOF OSEM (40it, 1ss) without scatter estimation.
+Simulation and reconstruction were then performed in the following three ways:
   1. 1.
-    The full dataset (556 million events) was reconstructed using OSEM (4it, 20ss) using the EM algorithm
+    A 9 minute acquisition (high count) was simulated in GATE that included all relevant physical phenomenon in a typical PET scan.
+The data was then reconstructed using LM-TOF OSEM (80it, 1ss) with attenuation, random, and scatter modeling.
+PSF modeling was implemented via 3 dimensional blurring in object space using a Gaussian kernel with a full width half max of 4 mm.
   2. 2.
-    6.3% of the data set (35 million events) was reconstructed using OSEM (80it, 1ss) using the EM algorithm, representative of a low-count scenario.
+    The dataset above was subsampled into a 3 minute acquisition (low count), and an identical reconstruction was performed.
   3. 3.
-    6.3% of the dataset (35 million events) was reconstructed using the DIPRecon algorithm.
-The prior image $z$ used was the corresponding T1-weighted MRI image; the initial network prediction $\mathcal{N}(z|\theta_{0})$ was obtained via fitting to the previous low count reconstruction.
-The neural network was a UNet35 with similar architecture to the original DIP paper 8; more information on the network structure and reconstruction hyperparameters can be found in the publicly available code.
-III. Results
-The time required for reconstruction in the first SPECT use case was 412.8 s (3.44 s/iteration) for OSEM, 434.8 s (3.62 s/iteration) for BSREM, and 601.8 s (5.02 s/iteration) for KEM; sample reconstructions with corresponding RC curves are shown for each algorithm in Figure 2.
-BSREM clearly produced the least noisy images, and had the best RCs for the lungs, liver, kidneys, and bladder.
-KEM produced an image estimate with an intermediate level of noise, and had the best recovery coefficient for the salivary glands.
-OSEM yielded the noisiest images, and had the worst recovery coefficients in all regions except the salivary glands.
-Despite the high performance of BSREM, it was unable to accurately predict the regions of low uptake within the liver visible in the ground truth.
-Figure 2: Reconstructions of full body 177Lu-PSMA-617 Monte Carlo projection data; ground truth phantoms were obtained using the XCAT software and SPECT simulation was achieved using the SIMIND Monte Carlo program.
-Left: Central coronal slices corresponding to truth, OSEM (120it/8ss), BSREM (120it/8ss), and KEM (120it/8ss).
-Right: Recovery coefficient vs.
-iteration number for sample organs; line widths are representative of error estimates based on reconstructions of 5 independent noise realizations.
-Sample reconstructions and RC curves for the SPECT motion correction use case are shown in Figure 3.
-The lesions appear sharper when motion correction is used; motion correction applied during reconstruction via Equation 3 produced lesions of higher uptake than when correction was applied post-reconstruction.
-This is quantitatively observed in the RC curves of Figure 3, where motion correction applied during reconstruction tends to produce the highest RCs throughout all iterations.
-After 100 iterations, the RCs were larger when MC was used during reconstruction than when it was used post reconstruction ($p<10^{-6}$).
-Figure 3: Reconstructions of full body 177Lu-PSMA-617 Monte Carlo projection data with breathing motion simulated; ground truth phantoms were obtained using the XCAT software and SPECT simulation was achieved using the SIMIND Monte Carlo program.
-Reconstruction was performed with OSEM (8ss) up to 100 iterations using attenuation/PSF modeling and TEW scatter correction; images shown correspond to reconstruction after 20 iterations.
-From left to right: ground truth phantom, reconstructed image without MC, reconstructed image with MC applied after reconstruction, reconstructed image with MC applied during reconstruction, and lesion recovery coefficients as a function of iteration number.
-Green boxes are drawn around the lesion in the liver.
-The time required to reconstruct the clinical SPECT patient data was 464.7s (9.29 s/iteration) for OSEM, 453.1s (9.06 s/iteration) for BSREM, and 1825s (36.50 s/iteration) for KEM.
-Central coronal slices are shown in Figure 4; the patient had a region of high uptake value in the liver, indicative of a liver lesion.
-BSREM and KEM appear to confine activity within the boundary of the kidneys, with BSREM exhibiting slightly more of this behaviour.
-In addition, OSEM produced the most observed noise in the liver.
-A common artifact of both BSREM and KEM, however, is the noise in the liver lesion; use of anatomical nearest neighbours may cause the lesion activity noise profile to resemble the CT noise profile.
-The mean predicted counts in the kidneys were 1.98 (OSEM), 2.07 (BSREM), and 2.02 (KEM); these relative proportions are consistent with the RCs of Figure 2, suggesting that BSREM may yield the most accurate recovery coefficients for the kidneys.
-Figure 4: Coronal slices of activity concentration (colored) and Hounsfield Unit values (greyscale) reconstructed using OSEM (50it/8ss), BSREM (50it/8ss), and KEM (50it/8ss).
-Data corresponds to the first time point of patient 4 from the “Lu-177 DOTATATE Anonymized Patient Datasets: Multi-Time Point Lu-177 SPECT/CT Scans” dataset31.
-Reconstructing the LM TOF PET GATE output data required multiple steps.
-While initial processing of the ROOT listmode data is time consuming, PyTomography can save this data in an accessible format for reuse.
-  • Opening the ROOT list mode data ($\sim$ 50 GB) of the normalization scan and computation of the normalization factor $\eta$ (approximately 50 minutes).
-When the normalization factor is saved, it requires approximately 5 s to read.
-  • Opening the ROOT list mode data ($\sim$ 74 GB) of the phantom scan (approximately 1 hour). When saved in an accessible format, reading requires approximately 5 s.
-  • Initial configuration of the system matrix ($7.8$ s).
-  • Reconstruction of the PET listmode data.
-The high count scenario using OSEM (4it 20ss) required 185.6 s (46.4 s per iteration), the low count scenario using OSEM (80it 1ss) required 201.0 (2.5 s per iteration), and the low count scenario using DIPRecon required 3135.6 s (31.3 s per iteration)
-Corresponding phantom reconstructions are shown in Figure 5; bias and noise in grey and white matter is shown as a function of iteration number in Figure 6.
-Due to finite resolution of the imaging system causing activity “leakage”, activity quantification in the grey matter region was consistently underestimated while activity quantification in the white matter region was consistently overestimated.
-The DIPRecon on the low count data had similar noise levels in grey matter to OSEM on the high count data, demonstrated both qualitatively and quantitatively in each Figures 5 and 6 respectively.
-Figure 5: PET reconstruction of GATE simulated data corresponding to a ultra-high resolution 18F-FDG PET/ T1-weighted MRI brain image34.
+    A low count reconstruction was performed without PSF modeling and used in the DIPRecon algorithm; no PSF modeling was used because (i) the DIP performs better on images with high frequency noise, and (ii) the DIPRecon algorithm already has noise reduction properties.
+The prior image $z$ used was the corresponding T1-weighted MRI image; the initial network prediction $\mathcal{N}(z|\theta_{0})$ was obtained via fitting to the low count reconstruction.
+The neural network was a UNet [27] with similar architecture to the original DIP paper [8]; more information on the network structure and reconstruction hyperparameters can be found in the publicly available code.
+The time required for the complete TOF-SSS estimation was 665 s (high count) and 577.5 s (low count), the time required for the OSEM reconstructions were 3.31 s/it (high count) and 1.18 s/it (low count), and the time required for the deep image prior reconstruction was 289.1 s (initial fit) and 21.53 s/it (reconstruction).
+Corresponding phantom reconstructions, as well as bias and noise in grey and white matter as a function of iteration number, are shown in Figure 4.
+Figure 4: Left: PET reconstruction of GATE simulated data corresponding to a ultra-high resolution 18F-FDG PET/ T1-weighted MRI brain image [25].
 From top to bottom: axial, coronal, and sagittal slices.
-From left to right: ground truth MRI (non contrast adjusted), ground truth PET uptake, high count OSEM reconstruction (20it/4ss), low count OSEM reconstruction (80it/1ss), and low count DIP reconstruction.
-Figure 6: Variation in absolute bias and standard deviation throughout iterative reconstruction corresponding to reconstructed images of Figure 5.
-Subplots correspond to the grey matter region (left) and white matter region (right).
+From left to right: ground truth MRI (non contrast adjusted), ground truth PET uptake, high count OSEM reconstruction (80it/1ss), low count OSEM reconstruction (80it/1ss), and low count DIP reconstruction.
+Right: Variation in absolute bias and standard deviation throughout iterative reconstruction.
+Subplots correspond to the grey matter region (upper) and white matter region (lower).
 For the low count rate scenario, points are plotted every 10 iterations.
 For the low count rate scenario with OSEM, the first point corresponds to the 10th iteration.
-The bias was negative in grey matter and positive in white matter.
-IV. Discussion
-This technical report presented the software architecture of PyTomography, and showed use cases on both SPECT and PET data.
-The short times required for reconstruction using this tool can enable researchers to perform extensive studies on phantom and patient data that may include multiple patients, noise realizations, and reconstruction algorithms.
-Validation of PyTomography against alternative reconstruction libraries is shown in the appendix.
-Example applications in 177Lu-SPECT imaging, which are relevant for currently FDA approved radiopharmaceutical therapies, explored the use of the OSEM, BSREM, and KEM, reconstruction algorithms.
-It should be emphasized that the purpose of this technical report was not to rank the different algorithms, but rather to demonstrate the capabilities of the software library.
-While some evidence was shown to favour BSREM over OSEM and KEM for kidney and liver dosimetry based on better recovery coefficients (Figure 2), a truly comprehensive study that aims to quantitatively compare algorithms should involve (i) multiple XCAT activity and anatomical configurations, (ii) inclusion of patient motion, (iii) effect of misalignment between SPECT and CT images, (iv) different hyperparameters in each reconstruction algorithm and (iv) should seek to use real patient data to validate any phantom studies.
-In addition, recovery coefficients alone are not sufficient to rank algorithms.
-Studies should also consider source-to-background ratios as well as error analysis based on intra- and inter-patient variabilities, and clinical task performance.
-While the presented software framework enables such studies to be performed, it is out of scope for this publication.
-While the RDP-AP prior was featured heavily in this study, the framework of PyTomography is designed to foster the development and validation of other novel prior functions as well.
-In addition, while KEM was featured using anatomically-based basis functions, the library permits use of any external image for construction of the basis functions.
-Different priors and basis functions may also have significant implications for (i) quantitative PET/SPECT imaging in dosimetry, and (ii) qualitative observer based studies in lesion detection and classification.
-The motion correction example considered a simplified case of 4 discrete phases where motion correction transforms were obtained using ground truth attenuation maps.
-In practice, however, breathing motion is continuous, and motion correction transforms will need to be derived using alternative source of data.
-The scope of this technical report is merely to demonstrate that PyTomography contains the functionality for incorporating motion correction.
-Example applications in 18F-FDG PET imaging demonstrated the capabilities of LM reconstruction using the OSEM and DIPRecon algorithms.
-As demonstrated qualitatively in Figure 5 and quantitatively in Figure 6, reconstruction of the low count data using DIPRecon was able to reduce noise in the grey matter regions while preserving cortical structures.
-Once again, the purpose of this paper is merely to demonstrate the capabilities of PET LM reconstruction and use of the DIPRecon algorithm provided in the software library; extensive research on the validity of the DIPRecon method should seek to explore, for example, the effects of different neural network architectures, hyperparameters, and misalignment between the PET/MRI images.
-While this paper has demonstrated implementations for parallel collimator SPECT and LM TOF PET, there exist many other modalities not presently included in the software, including (but not limited to) diverging/converging/pinhole SPECT, sinogram-based PET reconstruction, and various forms of computed tomography (CT) and magnetic resonance (MR) imaging.
+The bias was initially positive and became negative for all three cases in white matter.
+4 Impact
+In this work, we have motivated and presented the software architecture of PyTomography, and showed use cases on both SPECT and PET data.
+The short times required for reconstruction using this tool can enable researchers to perform extensive studies on digital/physical phantom and patient data that may include multiple patients, noise realizations, and reconstruction algorithms.
+To demonstrate (i) the impact of the software and (ii) the extent to which new research questions can be pursued, we will discuss each of the illustrative examples.
+A variety of algorithms were examined for 177Lu SPECT reconstruction; various quantitative metrics were estimated to highlight the differences between the available algorithms in the library.
+For both the simulated and clinical data, the algorithms yielding smallest to largest kidney concentration were BSREM-RDP, OSEM, KEM, and BSREM-RDPAP.
+Furthermore, the simulated data prediction from BSREM-RDPAP (807.02.17807.02.17807.02.17~{}807.02.17kBq/mL) was closest to the ground truth (843.36843.36843.36~{}843.36kBq/mL), suggesting that algorithms incorporating anatomical information may yield the most accurate kidney activity concentration.
+For both the simulated and clinical data, the order of algorithms yielding the lowest to highest noise were BSREM-RDP, BSREM-RDPAP, OSEM, and KEM; this is unsurprising as the RDP prior is designed to reduce noise.
+BSREM-RDPAP has a lower noise (6.9%) than OSEM (13.3%) but with higher uptake in the kidney (807.02 kBq/mL for BSREM-RDPAP vs.
+755.07 kBq/mL for OSEM) for the simulated data; BSREM-RDPAP thus has simultaneous improvements in both bias and noise reduction.
+For the bone lesion, algorithms incorporating anatomical information (BSREM-RDPAP and KEM) were able to contain activity mainly inside the bone; this is demonstrated in the axial slices shown on Figures 2 and 3.
+These algorithms also yielded higher bone lesion activity concentration predictions than OSEM for both the simulated/clinical data, with the simulated data concentrations (217.13217.13217.13~{}217.13kBq/mL for BSREM-RDPAP and 216.15216.15216.15~{}216.15kBq/mL for KEM) closer to the ground truth (819.79819.79819.79~{}819.79kBq/mL) than OSEM (198.88198.88198.88~{}198.88kBq/mL).
+For the liver lesion in the simulated data, however, use of the anatomical reconstruction algorithms decreased the activity concentration predictions (194.04194.04194.04~{}194.04kBq/mL for BSREM-RDPAP and 231.31231.31231.31~{}231.31kBq/mL for KEM) compared to OSEM (241.08241.08241.08~{}241.08kBq/mL).
+This occured because there is less anatomical difference in the surrounding tissue in the liver metastasis compared to the bone metastasis, so the smoothness prior merely blurred and reduced the activity concentration.
+These results, which are consistent with previous studies [6], suggest that anatomical priors may yield more accurate activity concentrations for bone lesion dosimetry.
+While this example focused primarily on organ and lesion dosimetry, PyTomography users will be to assess the merits of various reconstruction algorithms on any task specific metric.
+Example applications in 18F-FDG PET imaging demonstrated the capabilities of LM TOF SSS modeling and LM-TOF reconstruction using the OSEM and DIPRecon algorithms.
+As apparent in Figure 4, reconstruction of the low count data using DIPRecon was able to reduce noise in the grey matter regions while preserving cortical structures.
+While use of the DIPRecon on the low count data resulted in a slight increase in bias in the grey matter region, the corresponding noise bias curve demonstrates less noise than the high count reconstruction using OSEM.
+Extensive research on the validity of the DIPRecon method may seek to explore, for example, the effects of different neural network architectures, hyperparameters, and misalignment between the PET/MRI images.
+Since PyTomography directly interfaces with PyTorch, the process of designing and testing neural networks and optimization procedures for the DIPRecon method is straightforward.
+Users who wish to test different network/optimization configurations are encouraged to view the DIPRecon tutorial available on the documentation website.
+While this paper has demonstrated implementations for parallel collimator SPECT and LM-TOF PET, there exist many other modalities not presently included in the software, including (but not limited to) diverging/converging/pinhole SPECT, various forms of computed tomography, (CT), magnetic resonance (MR) imaging, and compton camera (CC) imaging.
 As displayed in Figure 1, projectors for different imaging modalities can be implemented using a variety of python based tools: it is hoped that the structure and flexibility of PyTomography will encourage contributions from various research groups interested in particular imaging systems.
+Implemented code does not necessarily need to be written directly in python; the parallelproj library used for PET modeling, for example, uses a CUDA C backend for its calculations.
+This flexibility permits contributions from users familiar with a variety of coding languages.
 Since the library is open source, newly developed functionality can be easily tested and verified by many independent research groups with their own data.
@@
   2. 2.
-    Incorporation of scatter and random estimation in PET reconstruction.
+    Expanded support for clinical PET data.
   3. 3.
-    Sinogram-based PET reconstructions.
+    Support for CT reconstruction
   4. 4.
-    Development of a 3D Slicer36 extension.
+    Development of a 3D Slicer [28] extension.
 In summary, we hope that we have made a compelling case for the benefit that this tool can bring to the medical imaging community.
 It can be used as a framework for studies that require fast reconstruction times, such as those that explore a wide range of reconstruction hyperparameters.
-It is well validated and may be used for accurate and reproducible quantitative imaging in both PET and SPECT; we particularly envision applications in the field of theranostics.
-A variety of developments will assist PyTomography in reaching the full potential of precision medicine.
-We thus highly encourage and welcome open-source contributions to assist in the items listed above; those who wish to contribute are encouraged to read the developers’ guide on the documentation website.
-V. Conclusion
+It may also be used by developers seeking to implement new system matrices and likelihoods for their own imaging systems.
+We highly encourage and welcome open-source contributions to assist in the items listed above; those who wish to contribute are encouraged to read the developers’ guide on the documentation website.
+5 Conclusions
 This work describes the python library PyTomography and highlights specific use cases in SPECT and PET imaging.
-The software architecture facilitates the development of different imaging modalities that can all interface with the same reconstruction algorithms.
-The purpose of PyTomography is to create a well-documented and computationally-efficient library for medical image reconstruction, where novel system modeling and reconstruction techniques are implemented, shared, and explored by experts in the community.
-VI. Appendix
-VI.A. Relevant Links
-Links to the various webpages of PyTomography are contained in Table 5: GitHub contains the source code, Readthedocs contains all relevant documentation, PyPI is the host of the built source files, and Discourse is a community forum for users and contributors.
-Table 5: Relevant Webpages
-Website Link
-Github https://github.com/qurit/PyTomography
-Readthedocs https://pytomography.readthedocs.io/en/latest/
-PyPI https://pypi.org/project/pytomography/
-Discourse https://pytomography.discourse.group/
-VI.B. Validation of PyTomography
-This section contains two examples used to validate PyTomography against other reconstruction software:
+The software architecture facilitates the development of different imaging modalities and likelihoods that can all interface with the same reconstruction algorithms.
+The goal of this research was to create a simple-to-use and computationally-efficient library for medical image reconstruction, where system modeling and reconstruction techniques are implemented, shared, and explored by experts in the community.
+Acknowledgements
+This work was supported by the Natural Sciences and Engineering Research Council of Canada (NSERC) CGS D Award 569711, NSERC Discovery Grants RGPIN-2019-06467 and RGPIN-2021-02965, as well as computational resources and services provided by Microsoft AI for Health.
+Appendix A Validation of PyTomography
+This section contains three examples used to validate PyTomography against other reconstruction software:
   1. 1.
     Monte Carlo SPECT Validation With STIR: Radioactivity concentrations in a digital phantom were selected to correspond to a 24-hour post injection prostate cancer patient (1700 MBq total activity) imaged via the prostate-specific membrane antigen (PSMA)-targeting radiopharmaceutical, 177Lu-PSMA-617.
-Simulated data was reconstructed with OSEM (2 iterations, 8 subsets) using (i) STIR and (ii) PyTomography.
+SPECT acquisition was simulated in SIMIND with dimensions $128\times 384$, a pixel size of $0.48~{}\text{cm}\times 0.48~{}\text{cm}$, 120 projections, an acquisition time of 15 s/projection, and use of Siemens medium energy collimators.
+Data were reconstructed with OSEM (2 iterations, 8 subsets) using (i) STIR and (ii) PyTomography.
 Coronal slices and 1D profiles of reconstructions are shown in Figure 5 and are nearly identical.
@@
   2. 2.
-    Clinical SPECT Validation With Vendor: A Jaszczak phantom was filled with Lu-177 and scanned at 180 projections ($128\times 128$ with a resolution of $4.8\text{mm}\times 4.8\text{mm}$) using a Siemens Symbia T Series SPECT/CT scanner with Medium Energy collimators.
-Reconstruction was performed with OSEM (4 iterations, 10 subsets) using (i) the manufacturer scanner software and (ii) PyTomography.
-Reconstructions and corresponding 1 dimensional profiles are shown in Figure 8; they appear similar.
-3D slicer was used to obtain sample segmentations of the hot and cold regions; the mean source-to-background ratios were 20.97 (vendor) and 21.76 (PyTomography).
-The image noise, quantified by the standard deviation of predicted counts in the cold region, were 0.217 counts (vendor) and 0.178 counts (PyTomography).
+    Clinical SPECT Validation With Siemens and MIM: A NEMA phantom was filled with 177Lu at a 10:1 source to background ratio and scanned at 96 projections ($128\times 128$ with a resolution of $4.8\text{mm}\times 4.8\text{mm}$) for 15 s per projection using a Siemens Symbia T Series SPECT/CT scanner with Medium Energy collimators.
+Reconstruction was performed with OSEM (4 iterations, 8 subsets) with attenuation/PSF modeling and TEW scatter correction, using (i) the manufacturers software: Siemens Symbia Intevo Excel v9.0.31.3, (ii) MIM SPECTRA v7.2.1, and (iii) PyTomography.
+Reconstructions and corresponding 1 dimensional profiles are shown in Figure 6; the PyTomography and Siemens reconstructions are qualitatively more similar than the MIM reconstruction.
+3D slicer was used to obtain sample segmentations of (i) the largest NEMA sphere, (ii) the central cold portion, and (iii) the warm background region.
+The mean counts in the largest NEMA sphere were 14.07 (Siemens), 12.96 (MIM), and 13.49 (PyTomography).
+The mean counts in the central cold portion were 0.732 (Siemens), 0.704 (MIM), and 0.688 (PyTomography).
+The image noise, quantified by the standard deviation of predicted counts in the warm region, were 0.196 (Siemens), 0.168 (MIM), and 0.179 (PyTomography).
   3. 3.
-    Monte Carlo PET Validation With CASToR: The ultra-high resolution 18F-FDG PET/ T1-weighted MRI brain image34 used in the PET example (556 million events) was reconstructed using LM TOF OSEM (2 iterations, 21 subsets) with both PyTomography and CASToR.
+    Monte Carlo PET Validation With CASToR: The ultra-high resolution 18F-FDG PET/ T1-weighted MRI brain image [25] used in the PET example was simulated in GATE without random, scatter, attenuation, or positron range modeling to obtain a ultra high count dataset consisting only of primary events.
+The data were reconstructed using LM TOF OSEM (2 iterations, 21 subsets) with both PyTomography and CASToR.
 Resulting reconstructions are shown in Figure 7.
@@
 The noise in the image was estimated as the standard deviation of counts in white matter divided by the mean activity in white matter; the values were 0.329 (PyTomography) and 0.321 (CASToR).
-Apart from processing the ROOT files and building the sensitivity image, the time required for reconstruction in PyTomography was 14 s to build the system matrix and required 59 s per iteration for reconstruction.
+Apart from processing the ROOT files and building the sensitivity image, the time required for reconstruction in PyTomography was 14 s to build the system matrix and required 59 s/it for reconstruction.
 CASToR reconstruction was ran on a separate HP Z8 Computer Desktop with 2 Intel Xeon Gold 6136 CPUs at 3.00 GHz with a total of 48 CPU cores.
@@
 1D profiles correspond to the horizontal lines shown on the coronal slices.
-Figure 8: Sample axial slices (top) and 1D profiles (bottom) of a 177Lu Jaszczak phantom reconstructed using OSEM (4it/10ss) with attenuation/PSF modeling and triple energy scatter correction.
-Reconstructions were performed using (i) the scanner software, and (ii) PyTomography.
+Figure 6: Sample axial slices (top) and 1D profiles (bottom) of a 177Lu NEMA phantom reconstructed using OSEM (4it/8ss) with attenuation/PSF modeling and triple energy scatter correction.
+Reconstructions were performed using (i) the Siemens scanner software, (ii) MIM SPECTRA Reconstruction, and (iii) PyTomography.
 1D profiles correspond to the horizontal lines shown on the axial slices.
@@
 1D profiles correspond to the horizontal lines shown on the coronal slices
-Acknowledgements
-This work was supported by the Natural Sciences and Engineering Research Council of Canada (NSERC) CGS D Award 569711, NSERC Discovery Grants RGPIN-2019-06467 and RGPIN-2021-02965, as well as computational resources and services provided by Microsoft AI for Health.

### v5

@@
 Abstract
-There is a need for open-source libraries in emission tomography that (i) use modern and popular backend code to encourage community contributions and (ii) offer support for the multitude of reconstruction algorithms available in recent literature, such as those that employ artificial intelligence.
+There is a need for open-source libraries in emission tomography that (i) use modern and popular backend code to encourage community contributions and (ii) offer support for the multitude of reconstruction techniques available in recent literature, such as those that employ artificial intelligence.
 The purpose of this research was to create and evaluate a GPU-accelerated, open-source, and user-friendly image reconstruction library, designed to serve as a central platform for the development, validation, and deployment of various tomographic reconstruction algorithms.
@@
 Nr. Code metadata description Please fill in this column
-C1 Current code version v3.0.0
+C1 Current code version v3.2.3
 C2 Permanent link to code/repository used for this code version https://github.com/qurit/PyTomography
@@
 C6 Software code languages, tools, and services used python
-C7 Compilation requirements, operating environments & dependencies python >=>=> = 3.11
+[…]
 C8 If available Link to developer documentation/manual https://pytomography.readthedocs.io/en/latest/
@@
 Valuable past efforts have focused on CUDA programming in languages such as C/C++ to accelerate reconstruction times in SPECT [16] and PET [16, 17] reconstruction through the use of multi-threading.
-While some contributors to an open-source framework may prefer to use C/C++ programming, others may be unfamiliar with these languages; any tool that prioritizes community contributions should permit both options.
-Some of the advantages of a python framework are that (i) it is a widely known programming language and thus encourages community involvement and (ii) it permits the embedding of C/C++ code for particular applications.
-Python can directly utilize CUDA acceleration via the PyTorch library, for example, allowing users unfamiliar with C/C++ code to achieve GPU acceleration.
-To this end, we developed the python library PyTomography with the priorities to: (i) implement standard and traditional imaging modalities and reconstruction algorithms, (ii) implement and disseminate of recent research developments, such as the deep image prior [8], and (iii) encourage community involvement via extensive documentation and user-friendly tutorials.
-Based on the motivations previously described, the backend of library has been developed using PyTorch because it maintains a similar functional structure to the widely used NumPy library.
-Other libraries are used for more particular tasks; the python library paralleproj [19], which was developed with CUDA C backend code, is used for line integral computation in PET projectors.
-In general, code written in any language can be integrated in PyTomography provided a python interface is available.
+While some contributors to an open-source framework may be comfortable with C/C++ programming, others may be unfamiliar with these languages; any tool that prioritizes community contributions should be designed with a front end application programming interface (API) that permits either of these options.
+Some of the advantages of a python front end API are that (i) it is a widely known programming language and thus facilitates community use, (ii) it enables direct interaction with other python libraries such as PyTorch for deep learning, and (iii) it enables backend development and contributions from other programming languages if necessary.
+This may address limitations of the C/C++ front end libraries STIR and CASToR, which are currently the most popular open-source reconstruction frameworks in emission tomography.
+To this end, we developed the python library PyTomography with the priorities to: (i) implement standard and traditional imaging modalities and reconstruction algorithms, (ii) disseminate recent research developments, such as the deep image prior [8], and (iii) encourage community involvement via extensive documentation and user-friendly tutorials.
+Based on the motivations previously described, the main functionality of the library has been developed using PyTorch.
+Other libraries are used for more particular tasks; the python library paralleproj [19] is used for line integral computation in PET projectors.
 While the present focus of PyTomography is the development and validation of reconstruction algorithms for SPECT and PET, implementation of other imaging modalities and reconstruction algorithms can be added using the building blocks provided.
-Overall, the aim of this study is to introduce the community to the PyTomography package, demonstrate its current capabilities and flexibility, and to encourage community involvement in the continuation of its development.
-To demonstrate the versatility of the software, example applications are explored in (i) SPECT reconstruction using PyTorch based rotate+sum projectors, and (ii) listmode (LM) time-of-flight (TOF) PET reconstruction.
-In what follows, we elaborate on our methods and results for PyTomography, which is made publicly available to download.
+In addition to the examples provided in this paper, PyTomography has support for high energy SPECT reconstruction via interface with the SPECTPSFToolbox [20], reconstruction using multiple SPECT photopeaks simultaneously, and region-based uncertainty estimation [21]; such features are not available in the other open-source libraries.
+There is ongoing work in the development of an associated graphical user interface extension in 3D slicer for SPECT reconstruction [22]; the corresponding repository can be found on the PyTomography Github organization page.
+Overall, the aim of this study is to introduce the community to the PyTomography software project, demonstrate its current capabilities and flexibility, and to encourage community involvement in the continuation of its development.
+Example applications are explored in SPECT and PET reconstruction, though the library also has capabilities for clinical CT reconstruction as well.
+In what follows, we elaborate on our methods and results for PyTomography.
 2 Software description
@@
 When $A1$ or $A^{T}1$ is written, “1” is to be interpreted as a vector (in the appropriate vector space) with all elements equal to 1.
+When classes/functions are quoted, they refer to the either (i) functionality of the current PyTomography version or (ii) functionality of dependencies (see pyproject.toml file) of the current PyTomography version.
 2.1.1 Projectors
@@
 In emission tomography, analytical modeling of the data expectation $\bar{g}$ can be achieved via $\bar{g}=Hf+s$ where $s$ is an additive term containing estimates for scattered events (SPECT) or scatter and random coincidences (PET).
-In CT, analytical modeling is achieved via $\bar{g}=be^{-Hf}$ where $b$ denotes the rate of emissions from the transmission source to a particular measurement.
 The operation of $H$ on an object vector is known as forward projection (FP), while the operation of $H^{T}$ on a projection vector is known as back projection (BP).
@@
     $H=\sum_{n}v_{n}\otimes\left(\mathcal{T}_{\text{fov}}^{(n)}~{}\mathcal{P}_{x}~{}\mathcal{T}_{\text{cdr}}^{(n)}\mathcal{T}_{\text{att}}^{(n)}~{}\mathcal{T}_{\text{trans}}^{(n)}~{}\mathcal{T}_{\text{rot}}^{(n)}\right)$ (1)
-    where $n$ designates the projection angle, $\mathcal{T}_{\text{rot}}^{(n)}$ implements a rotation, $\mathcal{T}_{\text{trans}}^{(n)}$ implements a translation that accounts for non-central projections, $\mathcal{T}_{\text{att}}^{(n)}$ implements the adjustment for photon attenuation, $\mathcal{T}_{\text{cdr}}^{(n)}$ implements depth-dependent collimator blurring, $\mathcal{P}_{x}$ is summation along the $+x$ axis, and $\mathcal{T}_{\text{fov}}^{(n)}$ applies a mask in projection space that accounts for a finite field of view, $v_{n}$ is a unit vector of length $N$ with 1 in component $n$, and $\otimes$ is the outer product.
+    where $n$ designates the projection angle, $\mathcal{T}_{\text{rot}}^{(n)}$ implements a rotation, $\mathcal{T}_{\text{trans}}^{(n)}$ implements a translation that accounts for non-central projections, $\mathcal{T}_{\text{att}}^{(n)}$ implements the adjustment for photon attenuation, $\mathcal{T}_{\text{cdr}}^{(n)}$ implements depth-dependent collimator blurring, $\mathcal{P}_{x}$ is summation along the $+x$ axis, $\mathcal{T}_{\text{fov}}^{(n)}$ applies a mask in projection space that accounts for a finite field of view, $v_{n}$ is a unit vector of length $N$ with 1 in component $n$, and $\otimes$ is the outer product.
+The most computationally expensive operations in this sequence are $\mathcal{T}_{\text{cdr}}^{(n)}$ and $\mathcal{T}_{\text{rot}}^{(n)}$, which are implemented using the PyTorch functions rotate (from torchvision.transforms.functional) and conv1d/conv2d (from torch.nn).
   • 2.
@@
     where $\mathcal{T}_{rr}$ implements object-space resolution modeling (typically via a uniform Gaussian kernel), $\mathcal{P}_{\Sigma}$ implements line integral computation (listmode/sinogram), $\mathcal{T}_{\eta}$ adjusts for detector normalization, and $\mathcal{T}_{\omega}$ adjusts for attenuation effects.
-PyTomography can be used to construct projectors that extend projection space; such functionality is necessary, for example, when modeling multiple photopeaks in SPECT or when using motion correction.
+The most computationally expensive operation is $\mathcal{P}_{\Sigma}$, which is implemented using joseph3d_fwd (non-TOF) and joseph3d_fwd_tof_sino/ joseph3d_fwd_tof_lm (TOF) from parallelproj.
+PyTomography can be used to construct projectors that extend projection space using the ExtendedSystemMatrix class; such functionality is necessary, for example, when modeling multiple photopeaks in SPECT or when using motion correction.
 In such situations, the extended projection space has dimensionality $(N,...)$ where ...
@@
   • 2.
-    Multiple SPECT Photopeaks: In this scenario, $A_{n}$ are scaling factors that adjust for detector sensitivities and branching ratios, $B_{n}$ are identity mappings, and $H_{n}$ represents the system matrix corresponding to a particular photopeak.
+    Multiple SPECT Photopeaks: In this scenario, $A_{n}$ are scaling factors that adjust for detector sensitivities and branching ratios between different emission peaks, $B_{n}$ are identity mappings, and $H_{n}$ represents the system matrix corresponding to a particular photopeak.
 The extended projection space has dimensions $(N,...)$ where $N$ is the number of different photopeaks.
@@
 In PyTomography, the “likelihood” $L(g|f,H)$ refers to an objective function to be maximized; as such, the “likelihood” $L(g|f,H)$ is often used interchangeably with the statistical “log-likelihood” $\log\tilde{L}(g|f,H)$ for a particular modality.
-Likelihood functions for different modalities are shown in Table 2.
-Table 2: Likelihoods and corresponding gradients for various medical imaging modalities; $s$ represents a scatter estimate used in SPECT/PET while $b$ represents a scaling factor used in CT.
-The CT likelihood corresponds to an ideal scenario where the number of counts in each sinogram bin follows a Poisson distribution; this is not necessarily the case in real systems with complex electronic noise.
-Likelihood $L(g|f,H)$ $\nabla_{f}L(g|f,H)$
-SPECT/PET $\sum_{i}\left(g_{i}\log\left[Hf+s\right]_{i}-\left[Hf+s\right]_{i}\right)$ $H^{T}\left(\frac{g}{Hf+s}\right)-H^{T}1$
-CT $\sum_{i}\left(g_{i}\log\left[be^{-Hf}\right]_{i}-\left[be^{-Hf}\right]_{i}\right)$ $H^{T}(be^{-Hf}-g)$
-Weighted MSE $-\frac{1}{2}\sum_{i}\left[W\left(g-Hf\right)^{2}\right]_{i}$ $H^{T}W(g-Hf)$
+Likelihood functions for different modalities are shown in Table 2; evaluation of a likelihood and its derivatives require forward and back projection of system matrices.
 2.1.3 Reconstruction Algorithms
@@
 where the ellipsis “...” includes all additional hyperparameters required for the algorithm; the ellipsis will be assumed implicit for the remainder of this paper.
-The main algorithms in PyTomography are preconditioned gradient ascent (PGA) algorithms; these algorithms aim to maximimize the likelihood via iterative updates that take the form
+The main algorithms in PyTomography are preconditioned gradient ascent (PGA) algorithms.
+These algorithms aim to maximimize the likelihood via iterative updates that take the form
 $f^{n+1}=f^{n}+C^{n}(f^{n},H^{n})\left[\nabla_{f}L(g^{n}|f^{n},H^{n})-\beta^{n}\nabla_{f}V(f^{n})\right]$ (5)
@@
 Various PGA algorithms available in PyTomography are shown in Table 2; they are characterized by their preconditioner factor.
-Certain algorithm names in the literature correspond to specific realizations of the generic algorithms in Table 3 with particular likelihoods in Table 2.
-The OS-SART algorithm, for example, corresponds to the use of OSEM with the weighted MSE likelihood (with weights $W=1/(H1)$).
-Table 3: Sample algorithms available in PyTomography: shown are ordered subset expectation maximum (OSEM), ordered subset maximum a posteriori one step late (OSMAPOSL), block sequential regualized expectation maximum (BSREM), and kernelized expectation maximum (KEM).
-$n$ is used as a superscript (not an exponent) for all variables.
-For BSREM, $\alpha^{n}$ is a relaxation parameter.
-Algorithm Preconditioner factor $C^{n}(f^{n},H^{n})$
-OSEM [20] $f^{n}/\left(\left(H^{n}\right)^{T}1\right)$ and $V(f^{n})=0$
-OSMAPOSL [21] $f^{n}/\left(\left(H^{n}\right)^{T}1+\nabla_{f}V(f^{n})\right)$
-BSREM [22] $f^{n}\alpha^{n}/\left(c^{n}H^{T}1\right)$
 Prior functions $V(f)$ used in reconstruction often take the form
@@
 When $w_{rs}$ is non-zero only for neighbouring voxels, the prior is known as a nearest neighbour prior; various nearest neighbour priors implemented in PyTomography are shown in Table 2.
-While the weighting factor $w_{rs}$ is typically proportional to inverse Euclidean distance, it is also possible to allow it to depend on an additional anatomical image $\chi$ whereby $w_{rs}=w_{rs}(\chi_{r},\chi_{s})$ Since anatomical information is used, this is commonly referred to as an anatomical prior (AP); variations of this technique have recently become popular in clinical practice due to enhanced lesion quantitation and detectability in bone SPECT/CT [6].
+While the weighting factor $w_{rs}$ is typically proportional to inverse Euclidean distance, it may also be dependent on an additional anatomical image $\chi$ whereby $w_{rs}=w_{rs}(\chi_{r},\chi_{s})$ Since anatomical information is used, this is commonly referred to as an anatomical prior (AP); variations of this technique have recently become popular in clinical practice due to enhanced lesion quantitation and detectability in bone SPECT/CT [6].
 In this paper, use of an AP refers to using only the 8 most similar neighbours based on absolute differences of HU values in a CT scan or attenuation map.
-Table 4: Sample nearest neighbour functions available in PyTomography; these priors are characterized by their image weighting function $\phi(f_{r},f_{s})$.
-Prior $\phi(f_{r},f_{s})$
-Quadratic $\left((f_{r}-f_{s})/\delta\right)^{2}/4$
-Log-cosh $\tanh((f_{r}-f_{s})/\delta)$
-Relative difference penalty (RDP) [23] $\left((f_{r}-f_{s})^{2}\right)/\left(f_{r}+f_{s}+\gamma|f_{r}-f_{s}|\right)$
+The prior used in subsequent examples in this paper is the relative difference penalty (RDP) [23], since it is commonly used in vendor reconstruction.
+Table 2: Sample PyTomography classes and their corresponding identifying features.
+$s$ corresponds to additive term in likelihoods.
+The sample algorithms shown are the ordered subset expectation maximum (OSEM) [24], ordered subset maximum a posteriori one step late (OSMAPOSL) [25], block sequential regularized expectation maximum (BSREM) [26].
+$n$ is used as a superscript (not an exponent) for all variables.
+For BSREM, $\alpha^{n}$ is a relaxation parameter.
+The sample nearest neighbour prior functions are characterized by their image weighting function $\phi(f_{r},f_{s})$.
+Projector classes System matrix $H$
+SPECTSystemMatrix Equation 1
+PETLMSystemMatrix Equation 2
+Likelihood classes Gradient $\nabla_{f}L(g|f,H)$
+PoissonLogLikelihood $H^{T}\left(\frac{g}{Hf+s}\right)-H^{T}1$
+NegativeMSELikelihood $H^{T}W(g-(Hf+s))$
+Algorithm classes Preconditioner factor $C^{n}(f^{n},H^{n})$
+OSEM $f^{n}/\left(\left(H^{n}\right)^{T}1\right)$ and $V(f^{n})=0$
+OSMAPOSL $f^{n}/\left(\left(H^{n}\right)^{T}1+\nabla_{f}V(f^{n})\right)$
+BSREM $f^{n}\alpha^{n}/\left(c^{n}H^{T}1\right)$
+Prior classes Weighting function $\phi(f_{r},f_{s})$
+QuadraticPrior $\left((f_{r}-f_{s})/\delta\right)^{2}/4$
+LogCoshPrior $\tanh((f_{r}-f_{s})/\delta)$
+RelativeDifferencePrior $\left((f_{r}-f_{s})^{2}\right)/\left(f_{r}+f_{s}+\gamma|f_{r}-f_{s}|\right)$
 PyTomography also offers support for the Deep Image Prior Reconstruction (DIPRecon) algorithm.
@@
 Metadata classes contain all auxiliary information needed for image reconstruction pertaining to a particular imaging system, such as voxel size, spatial dimensions, and information pertaining to the raw data of a particular modality.
+The input/output (I/O) module is used for reading various file formats; it currently supports SIMIND, DICOM (SPECT/CT), GATE (PET), and HDF5 listmode (PET) formats.
+Data reading consists of loading the generated files and constructing the appropriate data/metadata instances within the PyTomography framework.
+The I/O module is also used to read internal files of the library, such as coefficients to convert Hounsfield Units to attenuation coefficients, files containing SPECT collimator information, and files containing PET scanner geometry.
   2. 2.
@@
 While the input/output datatype to projectors is the PyTorch tensor class, internal code can utilize functionality from other libraries.
-Projector subclasses inherit from a generic “SystemMatrix“ class and implement a “forward”, “backward”, and “compute_normalization_factor” method.
-Projectors used in subset based algorithms should additionally implement “set_n_subsets”, “get_projection_subset” and ”get_weighting_subset”; the second method takes in projection data and returns subset data based on how the projector defines subsets, and the third gets the corresponding weight ($c^{n}$) for scaling preconditioners.
+Projector subclasses inherit from a generic SystemMatrix class and implement a forward, backward, and compute_normalization_factor method.
+Projectors used in subset based algorithms should additionally implement set_n_subsets, get_projection_subset and get_weighting_subset; the second method takes in projection data and returns subset data based on how the projector defines subsets, and the third gets the corresponding weight ($c^{n}$) for scaling preconditioners.
   3. 3.
@@
 Likelihoods $L(g|H,f)$ are initialized using a system matrix $H$, projection data $g$, and any additional dependencies; once initialized they act as functions of an image estimate $f$.
-Likelihoods inherit from a generic “Likelihood” class and should implement the “compute_gradient” method.
-Higher order derivatives of the likelihood may also be defined as seperate methods within the class.
+Likelihoods inherit from a generic Likelihood class and should implement the compute_gradient method.
+Higher order derivatives of the likelihood may also be defined as separate methods within the class.
   4. 4.
@@
 Statistical iterative algorithms require as input a likelihood instance, and may take as optional arguments prior functions and callbacks.
-Unlike projectors and likelihoods, algorithms do not inherit from a generic base class.
-In particular, however, linear PGAAs inherit from a generic “LinearPreconditionedGradientAscentAlgorithm” class and simply need to implement the “_linear_preconditioner_factor” and optionally reimplement the “__init__” method to take in required arguments.
-PyTomography also has an input/output (I/O) module for reading various file formats; it currently supports SIMIND, DICOM (SPECT/CT), GATE (PET), and HDF5 listmode (PET) formats.
-Data reading consists of loading the generated files and constructing the appropriate data/metadata instances within the PyTomography framework.
-The I/O module is also used to read internal files of the library, such as coefficients to convert Hounsfield Units to attenuation coefficients, files containing SPECT collimator information, and files containing PET scanner geometry.
-Novel development of a projector with associated data sources should thus consist of the following (i) implementation of metadata classes, (ii) I/O code to load data/metadata from the data sources, and (iii) development of a projector to interface with the data/metadata.
-Metadata class instances may be initialized by various data sources; for example, the same “SPECTProjMeta” class is initialized from SIMIND or DICOM I/O data loading functions, even though they represent different data sources.
-Metadata is thus linked to a modality and not to a specific data source.
-Deployment of the software was designed to maximize accessibility.
-In addition to placing the backend code on GitHub, the library was also uploaded to the Python Package Index (PyPi) so users can easily install the library using the simple “pip install pytomography” terminal command.
-A readthedocs documentation website was created that also contains a comprehensive set of tutorials in jupyter notebook format.
-While many of the tutorials are aimed at training end users, there also exist developer tutorials for implementing new projectors in the software.
-A discourse website was developed for two purposes: (i) to allow end users to ask questions about how to use the software and (ii) to permanently store and make public the questions/answers as they may be useful for other users of the library.
-Figure 1: Basic overview of the workflow of PyTomography when using statistical algorithms (left) with code representing a specific SPECT reconstruction use case (right).
-Shown are the minimum required inputs to the “__init__” function.
-the minimum required methods for each class, and sample subclasses available in the library.
-The “…” represents additional arguments that may be present in subclasses.
-In the SPECT use case, for example, the additional input argument “obj2obj_transforms” is implemented to take in operations that implement attenuation and PSF modeling.
-Not shown in the sample code: importing required libraries, loading the photopeak/scatter data, creating the attenuation/PSF transforms, and specifying the filepath “file_NM”.
+Linear PGAAs inherit from a generic LinearPreconditionedGradientAscentAlgorithm class and simply need to implement the _linear_preconditioner_factor and optionally re-implement the __init__ method to account for any additional required arguments.
+Sample code demonstrating the modularity of the library is shown in Figure 1 for a simplified SPECT use case (e.g.
+no attenuation, PSF, or scatter correction).
+Lines 1-9 demonstrate reading the data/metadata, lines 10-13 demonstrate constructing the projector using the loaded data/metadata, lines 14-17 demonstrate creating the likelihood using the system matrix and loaded data, and lines 18-22 demonstrate creating a reconstruction algorithm and obtaining a reconstructed image.
+A new contribution to the library might involve adding a new system model and replacing lines 1-13 while reusing lines 14-22.
+Another possible contribution might include a new reconstruction algorithm; in this case lines 18-22 would need to be replaced while the SPECT projector code can be reused.
+Figure 1: Sample code for a simplified SPECT reconstruction use case (e.g. no attenuation/PSF/scatter correction) that demonstrates the modular structure of the library.
 3 Illustrative examples
@@
 Validation against other software libraries is also shown in the Appendix, which contains (i) SIMIND simulated SPECT data reconstruction validated against STIR, (ii) DICOM SPECT data (obtained on a GE Tandem Discovery 670 Pro) validated against MIM’s SPECT reconstruction functionality, and (iii) GATE PET data reconstruction validated against CASToR.
-We chose to validate DICOM reconstruction against MIM, and not the Siemens internal reconstruction software, because it represents an external library with access to the same information as PyTomography.
+We chose to validate DICOM reconstruction against MIM, and not the GE internal reconstruction software, because it represents an external library with access to the same information as PyTomography.
 All computations were performed using a Microsoft Azure virtual machine (Standard NC6s v3) with a 6 CPUs (Intel(R) Xeon(R) CPU E5-2690 v4 @ 2.60GHz), 112 GB of RAM, and a TeslaV100 GPU.
@@
 This section highlights SPECT reconstructions using available reconstruction algorithms in PyTomography, demonstrating differences in quantitative metrics between different algorithms.
-Two examples are considered: (i) reconstruction of Monte Carlo SPECT data generated via SIMIND, and (ii) reconstruction of clinical SPECT data.
-In the clinical case, the patient received radiopharmaceutical therapy with 177Lu-PSMA for prostate cancer and was imaged 3 hours post injection; the XCAT phantom used as input for SIMIND was configured such that the kidney, liver, and lesion activity concentrations roughly approximated the clinical patient.
+Two examples are considered: (i) reconstruction of clinical SPECT data and (ii) reconstruction of Monte Carlo SPECT data generated via SIMIND.
+In the clinical case, the patient received radiopharmaceutical therapy with 177Lu-PSMA-617 for prostate cancer and was imaged 3 hours post injection; the XCAT phantom used as input for SIMIND was configured such that the kidney, liver, and lesion activity concentrations roughly approximated the clinical patient.
 In each case, the data correspond to acquisition on a GE Tandem Discovery 670 Pro SPECT system (MEGP collimators) with 60 projections (25 s per projection).
@@
 In each case, the system matrix employed attenuation and PSF modeling, and the likelihood function considered an additive scatter term obtained via the triple energy window technique.
-The reconstruction algorithms considered were OSEM (4it/6ss), BSREM-RDP (40it/6ss), BSREM-RDPAP (40it/6ss), and KEM (10it/6ss) [’it’ for iterations; and ’ss’ for subsets]; algorithms using RDP used the hyperparameters $\beta=0.3$, $\gamma=2$, RDPAP only considered the 8 nearest neighbours based on the attenuation map, and KEM used Eq.
+The reconstruction algorithms considered were OSEM (4it/6ss), BSREM-RDP (40it/6ss), BSREM-RDPAP (40it/6ss), and KEM (10it/6ss) [‘it’ for iterations; and ‘ss’ for subsets]; algorithms using RDP used the hyperparameters $\beta=0.3$, $\gamma=2$, RDPAP only considered the 8 nearest neighbours based on the attenuation map, and KEM used Eq.
 5 of Vuohijoki [6] for its filter.
-The time required for reconstruction of one bed position of the clinical data was 1.07 s/it (OSEM), 0.83 s/it (BSREM-RDP), 0.86 s/it (BSREM-RDPAP), and 1.07 s/it (KEM).
-The time required for reconstruction of the simulated data was 1.32 s/it (OSEM), 1.13 s/it (BSREM-RDP), 1.21 s/it (BSREM-RDPAP), and 1.52 s/it (KEM).
+The required computation times are shown in Table 3.
 For the clinical data, the left kidney was segmented using the CT-based TotalSegmentator model [27], and the bone metastasis in the sternum was segmented by first segmenting the lesion from the OSEM reconstruction, and then applying a threshhold mask greater than 0.150.150.15~{}0.15cm-1 so that only voxels within the bone were considered.
 A spherical mask was used in the liver for both the simulated and clinical data for measurement of noise in each reconstruction algorithm; noise was quantified as a percent by computing the standard deviation of uptake divided by the mean of uptake in that region.
-Reconstructions of the clinical and simulated data are shown in Figures 2 and 3 respectively.
-Figure 2: Top: Maximum intensity projections of reconstructed clinical SPECT data; the leftmost subplot corresponds to reconstruction using the MIM reconstruction software library.
-Three annotations are shown: (i) the left kidney, contained inside a pink box, with the corresponding activity concentration in kBq/mL (ii) the location of the spherical mask in the liver used for noise quantification with the corresponding noise in percent, and (iii) the location of the bone metastasis (blue arrow) with the corresponding activity concentration in kBq/mL.
-Bottom: Axial slice corresponding the hottest slice of the bone metastasis; SPECT uptake is shown in color and the attenuation map is shown in greyscale.
-Figure 3: Top: Maximum intensity projections of reconstructed simulated SPECT data; the leftmost subplot corresponds to the ground truth XCAT phantom used for simulation in SIMIND.
-Four annotations are shown: (i) the left kidney, contained inside a pink box, with the corresponding activity concentration in kBq/mL (ii) the location of the spherical mask in the liver used for noise quantification with the corresponding noise in percent, and (iii) the location of the bone metastasis (blue arrow) with the corresponding activity concentration in kBq/mL, and (iv) the location of the liver metastasis (purple arrow) with the corresponding activity concentration in kBq/mL.
-Bottom: Axial slice corresponding the hottest slice of the bone metastasis; SPECT uptake is shown in color and the attenuation map is shown in greyscale.
+Reconstructions of the clinical and simulated data are shown in Figure 2.
+Figure 2: DICOM SPECT reconstruction of clinical data (top) and SIMIND SPECT reconstruction of Monte Carlo data (bottom).
+Greyscale images correspond to maximum intensity coronal projections of reconstructed images.
+The annotations correspond to: (i) the left kidney, contained inside a pink box, with the corresponding activity concentration in kBq/mL (ii) the location of the spherical mask in the liver used for noise quantification with the corresponding noise in percent, (iii) the location of the bone metastasis (blue arrow) with the corresponding activity concentration in kBq/mL, and (iv) the location of the liver lesion (purple arrow) inserted in the SIMIND data.
+Axial slices are shown below each greyscale image that correspond to the slice containing the bone metastasis; the SPECT activity is shown with a magma colormap superimposed on a Greyscale CT image.
 3.2 PET
@@
   3. 3.
-    Use of the DIPRecon algorithm to enhance reconstruction
+    Use of the DIPRecon algorithm to enhance reconstruction; the DIPRecon class of PyTomography directly takes as input an instance of an nn.Module of PyTorch, highlighting the library capabilities to interface directly with neural networks.
 The simulated scanner consisted of $8$ rings of 338 mm radius with 32.25 mm ring spacing, and a TOF coincidence resolution of 550 ps; this fictional scanner is representative of a Siemens Biograph mMR with TOF capabilities.
@@
 The normalization $\eta$ was obtained via simulation of a thin cylindrical shell of radius 318318318~{}318mm that extended along the axial FOV of the scanner; normalization computation is available in PyTomography for basic polygonal PET geometries.
-In all cases, simulated data were reconstructed using a $128\times 128\times 96$ matrix size with voxel spacing of 222~{}2mm; 5 TOF bins were used with a total range of 30 cm.
+In all cases, simulated data were reconstructed using a $204\times 204\times 154$ matrix size with voxel spacing of 1.251.251.25~{}1.25mm; 5 TOF bins were used with a total range of 30 cm.
 Before reconstruction, the “randoms” contribution to each listmode event was estimated from delayed coincidences; this process consisted of (i) binning the delayed coincidence events into a sinogram (ii) smoothing the sinogram to reduce noise, and (iii) converting back into LM form for use in LM reconstruction.
@@
     A 9 minute acquisition (high count) was simulated in GATE that included all relevant physical phenomenon in a typical PET scan.
-The data was then reconstructed using LM-TOF OSEM (80it, 1ss) with attenuation, random, and scatter modeling.
-PSF modeling was implemented via 3 dimensional blurring in object space using a Gaussian kernel with a full width half max of 4 mm.
+The data was then reconstructed using LM-TOF OSEM (4it, 20ss) with attenuation, random, and scatter modeling.
+PSF modeling was implemented via 3 dimensional blurring in object space using a Gaussian kernel with a full width half max of 3 mm.
   2. 2.
-    The dataset above was subsampled into a 3 minute acquisition (low count), and an identical reconstruction was performed.
+    The dataset above was subsampled into a 3 minute acquisition (low count), and an identical reconstruction was performed except with 80it and 1ss.
   3. 3.
-    A low count reconstruction was performed without PSF modeling and used in the DIPRecon algorithm; no PSF modeling was used because (i) the DIP performs better on images with high frequency noise, and (ii) the DIPRecon algorithm already has noise reduction properties.
+    A low count reconstruction was performed without PSF modeling and used as an initial image for the DIPRecon algorithm; no PSF modeling was used in the initial image because (i) the DIP performs better on images with high frequency noise, and (ii) the DIPRecon algorithm already has noise reduction properties.
 The prior image $z$ used was the corresponding T1-weighted MRI image; the initial network prediction $\mathcal{N}(z|\theta_{0})$ was obtained via fitting to the low count reconstruction.
 The neural network was a UNet [30] with similar architecture to the original DIP paper [8]; more information on the network structure and reconstruction hyperparameters can be found in the publicly available code.
-The time required for the complete TOF-SSS estimation was 665 s (high count) and 577.5 s (low count), the time required for the OSEM reconstructions were 3.31 s/it (high count) and 1.18 s/it (low count), and the time required for the deep image prior reconstruction was 289.1 s (initial fit) and 21.53 s/it (reconstruction).
-Corresponding phantom reconstructions, as well as bias and noise in grey and white matter as a function of iteration number, are shown in Figure 4.
-Figure 4: Left: PET reconstruction of GATE simulated data corresponding to a ultra-high resolution 18F-FDG PET/ T1-weighted MRI brain image [25].
+Although PSF modeling was not used to obtain the initial image, it was employed during the iterative process of the DIPRecon algorithm.
+Corresponding phantom reconstructions, as well as bias and noise in grey and white matter as a function of iteration number, are shown in Figure 3. Required computation times are shown in Table 3.
+Figure 3: Top: PET reconstruction of GATE simulated data corresponding to a ultra-high resolution 18F-FDG PET/ T1-weighted MRI brain image [28].
 From top to bottom: axial, coronal, and sagittal slices.
-From left to right: ground truth MRI (non contrast adjusted), ground truth PET uptake, high count OSEM reconstruction (80it/1ss), low count OSEM reconstruction (80it/1ss), and low count DIP reconstruction.
-Right: Variation in absolute bias and standard deviation throughout iterative reconstruction.
-Subplots correspond to the grey matter region (upper) and white matter region (lower).
-For the low count rate scenario, points are plotted every 10 iterations.
-For the low count rate scenario with OSEM, the first point corresponds to the 10th iteration.
+From left to right: ground truth MRI (non contrast adjusted), ground truth PET uptake, high count OSEM reconstruction (4it/20ss), low count OSEM reconstruction (80it/1ss), and low count DIP reconstruction.
+Bottom: Variation in absolute bias and standard deviation throughout iterative reconstruction.
+Subplots correspond to the grey matter region (left) and white matter region (right).
+For the low count rate scenario, points are plotted every 10 iterations, and for the low count OSEM example, the first point corresponds to the 10th iteration.
 The bias was initially positive and became negative for all three cases in white matter.
+Table 3: Computational times for all use cases considered in this paper.
+Times required to open raw data are not considered.
+Times to construct SPECT system matrices and compute PET randoms in PyTomography are not considered since these are negligible.
+The notation “s / it” corresponds to the seconds per iteration for a given reconstruction algorithm, ‘SSS’ refers to the single scatter simulation in PET, and ‘SM’ refers to the creation of the system matrix.
+For the DIPRecon PET case, ‘LC’ refers to the prior low count reconstruction, since this is required to obtain an initial image for DIPRecon.
+SPECT DICOM Time Required
+OSEM (4it/6ss) (1.32 s/it/bed) $\times$ 4it $\times$ 2 bed $=$ 10.6 s
+BSREM-RDP (40it/6ss) (1.13 s/it/bed) $\times$ 40it $\times$ 2 bed $=$ 90.4 s
+BSREM-RDPAP (40it/6ss) (1.21 s/it/bed) $\times$ 40it $\times$ 2 bed $=$ 96.8 s
+KEM (10it/6ss) (1.52 s/it/bed) $\times$ 10it $\times$ 2 bed $=$ 30.4 s
+SPECT SIMIND Time Required
+OSEM (4it/6ss) (1.07 s/it) $\times$ 4it $=$ 4.3 s
+BSREM-RDP (40it/6ss) (0.83 s/it) $\times$ 40it $=$ 33.2 s
+BSREM-RDPAP (40it/6ss) (0.86 s/it) $\times$ 40it $=$ 34.4 s
+KEM (10it/6ss) (1.07 s/it) $\times$ 10it $=$ 10.7 s
+PET GATE Time Required
+High Count OSEM (4it/20ss) $(16.9~{}\text{s/it})&\times 4\text{it}+37.4~{}\text{s (SM)}\\ +&343.8~{}\text{s (SSS)}=\textbf{449.0~{}\text{s}}\end{aligned}$
+Low Count OSEM (80it/1ss) $(3.15~{}\text{s/it})&\times 80\text{it}+30.8~{}\text{s (SM)}\\ +&375.7~{}\text{s (SSS)}=\textbf{658.3~{}\text{s}}\end{aligned}$
+Low Count DIPRecon (100it) $(21.5~{}\text{s/it})&\times 100\text{it}+658.3~{}\text{s (LC)}\\ =&\textbf{2808.5~{}\text{s}}\end{aligned}$
+SPECT Validation Time Required
+PyTomo OSEM (2it/8ss) (2.70 s/it) $\times$ 2it $=$ 5.4 s
+STIR OSEM (2it/8ss) 51542.2 s
+PET Validation Time Required
+PyTomo OSEM (2it/21ss) (59 s/it) $\times$ 2it $=$ 118 s
+CASToR OSEM (2it/21ss) (455 s/it) $\times$ 2it $=$ 910 s
 4 Impact
-In this work, we have motivated and presented the software architecture of PyTomography, and showed use cases on both SPECT and PET data.
-The short times required for reconstruction using this tool can enable researchers to perform extensive studies on digital/physical phantom and patient data that may include multiple patients, noise realizations, and reconstruction algorithms.
-To demonstrate (i) the impact of the software and (ii) the extent to which new research questions can be pursued, we will discuss each of the illustrative examples.
+In this work, the software architecture of PyTomography was demonstrated and use cases were explored on both SPECT and PET data.
+The efficient reconstruction times, shown in Table 3, can enable researchers to perform extensive studies on digital/physical phantom and patient data that may include multiple patients, noise realizations, and reconstruction algorithms.
+A detailed discussion on each use case is included to highlight example analyses that can be performed with the software.
 A variety of algorithms were examined for 177Lu SPECT reconstruction; various quantitative metrics were estimated to highlight the differences between the available algorithms in the library.
 For both the simulated and clinical data, the algorithms yielding smallest to largest kidney concentration were BSREM-RDP, OSEM, KEM, and BSREM-RDPAP.
-Furthermore, the simulated data prediction from BSREM-RDPAP (807.02.17807.02.17807.02.17~{}807.02.17kBq/mL) was closest to the ground truth (843.36843.36843.36~{}843.36kBq/mL), suggesting that algorithms incorporating anatomical information may yield the most accurate kidney activity concentration.
+The simulated data prediction from BSREM-RDPAP (802.31802.31802.31~{}802.31kBq/mL) was closest to the ground truth (843.36843.36843.36~{}843.36kBq/mL), suggesting that algorithms incorporating anatomical information may yield the most accurate kidney activity concentration.
 For both the simulated and clinical data, the order of algorithms yielding the lowest to highest noise were BSREM-RDP, BSREM-RDPAP, OSEM, and KEM; this is unsurprising as the RDP prior is designed to reduce noise.
-BSREM-RDPAP has a lower noise (6.9%) than OSEM (13.3%) but with higher uptake in the kidney (807.02 kBq/mL for BSREM-RDPAP vs.
-755.07 kBq/mL for OSEM) for the simulated data; BSREM-RDPAP thus has simultaneous improvements in both bias and noise reduction.
-For the bone lesion, algorithms incorporating anatomical information (BSREM-RDPAP and KEM) were able to contain activity mainly inside the bone; this is demonstrated in the axial slices shown on Figures 2 and 3.
-These algorithms also yielded higher bone lesion activity concentration predictions than OSEM for both the simulated/clinical data, with the simulated data concentrations (217.13217.13217.13~{}217.13kBq/mL for BSREM-RDPAP and 216.15216.15216.15~{}216.15kBq/mL for KEM) closer to the ground truth (819.79819.79819.79~{}819.79kBq/mL) than OSEM (198.88198.88198.88~{}198.88kBq/mL).
-For the liver lesion in the simulated data, however, use of the anatomical reconstruction algorithms decreased the activity concentration predictions (194.04194.04194.04~{}194.04kBq/mL for BSREM-RDPAP and 231.31231.31231.31~{}231.31kBq/mL for KEM) compared to OSEM (241.08241.08241.08~{}241.08kBq/mL).
-This occured because there is less anatomical difference in the surrounding tissue in the liver metastasis compared to the bone metastasis, so the smoothness prior merely blurred and reduced the activity concentration.
+While BSREM-RDPAP has a lower noise than OSEM in each case, it also had higher uptake in the kidney: BSREM-RDPAP thus has simultaneous improvements in both bias and noise reduction.
+For the bone lesion, algorithms incorporating anatomical information (BSREM-RDPAP and KEM) were able to contain activity mainly inside the bone; this is demonstrated in the axial slices shown in Figure 2.
+These algorithms also yielded higher bone lesion activity concentration predictions than OSEM for both the simulated/clinical data, with the simulated data concentrations (347.77347.77347.77~{}347.77kBq/mL for BSREM-RDPAP and 220.56220.56220.56~{}220.56kBq/mL for KEM) closer to the ground truth (819.79819.79819.79~{}819.79kBq/mL) than OSEM (205.21205.21205.21~{}205.21kBq/mL).
+For the liver lesion in the simulated data, however, use of BSREM-RDPAP actually yielded a smaller activity concentration prediction (261.32261.32261.32~{}261.32kBq/mL) compared to OSEM (267.47267.47267.47~{}267.47kBq/mL).
+This occured because there was no anatomical difference in the surrounding tissue in the liver metastasis, while there was for the bone metastasis.
 These results, which are consistent with previous studies [6], suggest that anatomical priors may yield more accurate activity concentrations for bone lesion dosimetry.
@@
 While this paper has demonstrated implementations for parallel collimator SPECT and LM-TOF PET, there exist many other modalities not presently included in the software, including (but not limited to) diverging/converging/pinhole SPECT, various forms of computed tomography, (CT), magnetic resonance (MR) imaging, and compton camera (CC) imaging.
-As displayed in Figure 1, projectors for different imaging modalities can be implemented using a variety of python based tools: it is hoped that the structure and flexibility of PyTomography will encourage contributions from various research groups interested in particular imaging systems.
-Implemented code does not necessarily need to be written directly in python; the parallelproj library used for PET modeling, for example, uses a CUDA C backend for its calculations.
-This flexibility permits contributions from users familiar with a variety of coding languages.
-Since the library is open source, newly developed functionality can be easily tested and verified by many independent research groups with their own data.
+It is hoped that the modular structure and flexibility of PyTomography will encourage contributions from various research groups interested in particular imaging systems or reconstruction algorithms.
+Since the library is open source, newly developed functionality can be easily tested and verified by many independent research groups with their own private data.
 The immediate goals for future development in PyTomography at the time of publication include:
   1. 1.
-    Fast PSF modeling for high energy SPECT isotopes, such as 131I and 225Ac.
+    Expanded support for clinical PET data.
   2. 2.
-    Expanded support for clinical PET data.
+    Support for Monte Carlo scatter correction in SPECT reconstruction.
   3. 3.
@@
     Development of a 3D Slicer [22] extension.
-In summary, we hope that we have made a compelling case for the benefit that this tool can bring to the medical imaging community.
-It can be used as a framework for studies that require fast reconstruction times, such as those that explore a wide range of reconstruction hyperparameters.
-It may also be used by developers seeking to implement new system matrices and likelihoods for their own imaging systems.
-We highly encourage and welcome open-source contributions to assist in the items listed above; those who wish to contribute are encouraged to read the developers’ guide on the documentation website.
 5 Conclusions
@@
 Appendix A Validation of PyTomography
-This section contains three examples used to validate PyTomography against other reconstruction software:
+This section contains three examples used to validate PyTomography against other reconstruction software. Reconstruction times for cases 1 and 3 are shown in Table 3.
   1. 1.
@@
 The noise in the image was estimated as the standard deviation of counts in white matter divided by the mean activity in white matter; the values were 0.329 (PyTomography) and 0.321 (CASToR).
-Apart from processing the ROOT files and building the sensitivity image, the time required for reconstruction in PyTomography was 14 s to build the system matrix and required 59 s/it for reconstruction.
-CASToR reconstruction was ran on a separate HP Z8 Computer Desktop with 2 Intel Xeon Gold 6136 CPUs at 3.00 GHz with a total of 48 CPU cores.
-The reconstruction was performed with 24 CPU cores and reconstruction required approximately 455 s per full iteration.
+CASToR reconstruction was ran on a separate HP Z8 Computer Desktop with 2 Intel Xeon Gold 6136 CPUs at 3.00 GHz with a total of 48 CPU cores; this separate system was selected because it permitted use of more CPU cores in parallel during reconstruction, which is what CASToR is optimized for.
 The images shown in this section are displayed using the “nipy_spectral” colormap of matplotlib because it has significant contrast.

### v6

@@
 These algorithms aim to maximimize the likelihood via iterative updates that take the form
-$f^{n+1}=f^{n}+C^{n}(f^{n},H^{n})\left[\nabla_{f}L(g^{n}|f^{n},H^{n})-\beta^{n}\nabla_{f}V(f^{n})\right]$ (5)
-where $n$ is the subiteration index (used as a superscript and not an exponent for all variables), $f^{n}$ is the object estimate of subiteration $n$, $g^{n}$ specifies a subset of data used in subiteration $n$, $H^{n}$ is a system matrix projecting only to the current subset, $C^{n}(f^{n})$ is a preconditioner factor that depends on the object estimate, $V(f^{n})$ is a prior function used to regularize the algorithm, and $\beta^{n}=c^{n}\beta$ is a scaling factor for the prior function (scaled by the fraction of subset data $c^{n}$).
+$f^{(n+1)}=f^{(n)}+C^{(n)}(f^{(n)},H^{(n)})\left[\nabla_{f}L(g^{(n)}|f^{(n)},H^{(n)})-\beta^{(n)}\nabla_{f}V(f^{(n)})\right]$ (5)
+where $n$ is the subiteration index (used as a superscript and not an exponent for all variables), $f^{(n)}$ is the object estimate of subiteration $n$, $g^{(n)}$ specifies a subset of data used in subiteration $n$, $H^{(n)}$ is a system matrix projecting only to the current subset, $C^{(n)}(f^{(n)})$ is a preconditioner factor that depends on the object estimate, $V(f^{(n)})$ is a prior function used to regularize the algorithm, and $\beta^{(n)}=c^{(n)}\beta$ is a scaling factor for the prior function (scaled by the fraction of subset data $c^{(n)}$).
 These algorithms partition the data $g$ into $M$ approximately equal subsets; a single iteration of an algorithm is defined as the number of subiterations where all subsets have been used once.
@@
 $n$ is used as a superscript (not an exponent) for all variables.
-For BSREM, $\alpha^{n}$ is a relaxation parameter.
+For BSREM, $\alpha^{(n)}$ is a relaxation parameter.
 The sample nearest neighbour prior functions are characterized by their image weighting function $\phi(f_{r},f_{s})$.
@@
 NegativeMSELikelihood $H^{T}W(g-(Hf+s))$
-Algorithm classes Preconditioner factor $C^{n}(f^{n},H^{n})$
-OSEM $f^{n}/\left(\left(H^{n}\right)^{T}1\right)$ and $V(f^{n})=0$
-OSMAPOSL $f^{n}/\left(\left(H^{n}\right)^{T}1+\nabla_{f}V(f^{n})\right)$
-BSREM $f^{n}\alpha^{n}/\left(c^{n}H^{T}1\right)$
+Algorithm classes Preconditioner factor $C^{(n)}(f^{(n)},H^{(n)})$
+OSEM $f^{(n)}/\left(\left(H^{(n)}\right)^{T}1\right)$ and $V(f^{(n)})=0$
+OSMAPOSL $f^{(n)}/\left(\left(H^{(n)}\right)^{T}1+\nabla_{f}V(f^{(n)})\right)$
+BSREM $f^{(n)}\alpha^{(n)}/\left(c^{(n)}H^{T}1\right)$
 Prior classes Weighting function $\phi(f_{r},f_{s})$
@@
 Linear PGAAs inherit from a generic LinearPreconditionedGradientAscentAlgorithm class and simply need to implement the _linear_preconditioner_factor and optionally re-implement the __init__ method to account for any additional required arguments.
+Figure 1 shows a graphical depiction of the structure of the library.
 Sample code demonstrating the modularity of the library is shown in Figure 2 for a simplified SPECT use case (e.g.
@@
 Another possible contribution might include a new reconstruction algorithm; in this case lines 18-22 would need to be replaced while the SPECT projector code can be reused.
+Figure 1: Software architecture diagram for PyTomography.
+Metadata and projectors consist of an independent component of the library that can be used for analytical system modeling, but can also be used to construct likelihoods that can be passed to image reconstruction algorithms.
+Implementations of various components displayed here can be found in Table 2.
 Figure 2: Sample code for a simplified SPECT reconstruction use case (e.g. no attenuation/PSF/scatter correction) that demonstrates the modular structure of the library.
@@
 The normalization $\eta$ was obtained via simulation of a thin cylindrical shell of radius 318318318~{}318mm that extended along the axial FOV of the scanner; normalization computation is available in PyTomography for basic polygonal PET geometries.
-In all cases, simulated data were reconstructed using a $204\times 204\times 154$ matrix size with voxel spacing of 1.251.251.25~{}1.25mm; 5 TOF bins were used with a total range of 30 cm.
+In all cases, simulated data were reconstructed using a $204\times 204\times 154$ matrix size with voxel spacing of 1.251.251.25~{}1.25mm; 21 TOF bins were used with a total range of 30 cm.
 Before reconstruction, the “randoms” contribution to each listmode event was estimated from delayed coincidences; this process consisted of (i) binning the delayed coincidence events into a sinogram (ii) smoothing the sinogram to reduce noise, and (iii) converting back into LM form for use in LM reconstruction.
@@
     A 9 minute acquisition (high count) was simulated in GATE that included all relevant physical phenomenon in a typical PET scan.
-The data was then reconstructed using LM-TOF OSEM (4it, 20ss) with attenuation, random, and scatter modeling.
+The data was then reconstructed using LM-TOF OSEM for up to 27 iterations with 3 subsets with attenuation, random, and scatter modeling.
 PSF modeling was implemented via 3 dimensional blurring in object space using a Gaussian kernel with a full width half max of 3 mm.
   2. 2.
-    The dataset above was subsampled into a 3 minute acquisition (low count), and an identical reconstruction was performed except with 80it and 1ss.
+    The dataset above was subsampled into a 3 minute acquisition (low count), and an identical reconstruction was performed for up to 81 iterations with 1 subset.
   3. 3.
@@
 From top to bottom: axial, coronal, and sagittal slices.
-From left to right: ground truth MRI (non contrast adjusted), ground truth PET uptake, high count OSEM reconstruction (4it/20ss), low count OSEM reconstruction (80it/1ss), and low count DIP reconstruction.
-Bottom: Variation in absolute bias and standard deviation throughout iterative reconstruction.
+From left to right: ground truth MRI (non contrast adjusted), ground truth PET uptake, high count OSEM reconstruction (18it/3ss shown), low count OSEM reconstruction (54it/1ss shown), and low count DIP reconstruction.
+Bottom: Variation in absolute bias and standard deviation throughout iterative reconstruction; points are plotted at the iteration corresponding to the displayed images.
 Subplots correspond to the grey matter region (left) and white matter region (right).
-For the low count rate scenario, points are plotted every 10 iterations, and for the low count OSEM example, the first point corresponds to the 10th iteration.
 The bias was initially positive and became negative for all three cases in white matter.
@@
 The notation “s / it” corresponds to the seconds per iteration for a given reconstruction algorithm, ‘SSS’ refers to the single scatter simulation in PET, and ‘SM’ refers to the creation of the system matrix.
-For the DIPRecon PET case, ‘LC’ refers to the prior low count reconstruction, since this is required to obtain an initial image for DIPRecon.
-SPECT DICOM Time Required
+For the DIPRecon PET case, ‘IF’ refers to the initial model fit required for the DIPRecon.
+All timing performed on the same GPU-optimized system except for the CASToR PET validation example, which was ran on a CPU-optimized system.
+SPECT DICOM Computation Time
 OSEM (4it/6ss) (1.32 s/it/bed) $\times$ 4it $\times$ 2 bed $=$ 10.6 s
@@
 KEM (10it/6ss) (1.52 s/it/bed) $\times$ 10it $\times$ 2 bed $=$ 30.4 s
-SPECT SIMIND Time Required
+SPECT SIMIND Computation Time
 OSEM (4it/6ss) (1.07 s/it) $\times$ 4it $=$ 4.3 s
@@
 KEM (10it/6ss) (1.07 s/it) $\times$ 10it $=$ 10.7 s
-PET GATE Time Required
-High Count OSEM (4it/20ss) $(16.9~{}\text{s/it})&\times 4\text{it}+37.4~{}\text{s (SM)}\\ +&343.8~{}\text{s (SSS)}=\textbf{449.0~{}\text{s}}\end{aligned}$
-Low Count OSEM (80it/1ss) $(3.15~{}\text{s/it})&\times 80\text{it}+30.8~{}\text{s (SM)}\\ +&375.7~{}\text{s (SSS)}=\textbf{658.3~{}\text{s}}\end{aligned}$
-Low Count DIPRecon (100it) $(21.5~{}\text{s/it})&\times 100\text{it}+658.3~{}\text{s (LC)}\\ =&\textbf{2808.5~{}\text{s}}\end{aligned}$
-SPECT Validation Time Required
+PET GATE Computation Time
+High Count OSEM (27it/3ss) $(21.1~{}\text{s/it})&\times 27\text{it}+30.9~{}\text{s (SM)}\\ +&925.9~{}\text{s (SSS)}=\textbf{1525.5~{}\text{s}}\end{aligned}$
+Low Count OSEM (80it/1ss) $(6.69~{}\text{s/it})&\times 80\text{it}+30.8~{}\text{s (SM)}\\ +&1023.7~{}\text{s (SSS)}=\textbf{1589.8~{}\text{s}}\end{aligned}$
+Low Count DIPRecon (100it) $(25.5~{}\text{s/it})&\times 100\text{it}+663.8~{}\text{s (IF)}\\ =&\textbf{3209.6~{}\text{s}}\end{aligned}$
+SPECT Validation Computation Time
 PyTomo OSEM (2it/8ss) (2.70 s/it) $\times$ 2it $=$ 5.4 s
 STIR OSEM (2it/8ss) 51542.2 s
-PET Validation Time Required
+PET Validation Computation Time
 PyTomo OSEM (2it/21ss) (59 s/it) $\times$ 2it $=$ 118 s
-CASToR OSEM (2it/21ss) (455 s/it) $\times$ 2it $=$ 910 s
+CASToR OSEM 910 s
 4 Impact
@@
 For both the simulated and clinical data, the algorithms yielding smallest to largest kidney concentration were BSREM-RDP, OSEM, KEM, and BSREM-RDPAP.
-The simulated data prediction from BSREM-RDPAP (802.31802.31802.31~{}802.31kBq/mL) was closest to the ground truth (843.36843.36843.36~{}843.36kBq/mL), suggesting that algorithms incorporating anatomical information may yield the most accurate kidney activity concentration.
+The simulated data prediction from BSREM-RDPAP (807.88807.88807.88~{}807.88kBq/mL) was closest to the ground truth (843.36843.36843.36~{}843.36kBq/mL), suggesting that algorithms incorporating anatomical information may yield the most accurate kidney activity concentration.
 For both the simulated and clinical data, the order of algorithms yielding the lowest to highest noise were BSREM-RDP, BSREM-RDPAP, OSEM, and KEM; this is unsurprising as the RDP prior is designed to reduce noise.
@@
 For the bone lesion, algorithms incorporating anatomical information (BSREM-RDPAP and KEM) were able to contain activity mainly inside the bone; this is demonstrated in the axial slices shown in Figure 3.
-These algorithms also yielded higher bone lesion activity concentration predictions than OSEM for both the simulated/clinical data, with the simulated data concentrations (347.77347.77347.77~{}347.77kBq/mL for BSREM-RDPAP and 220.56220.56220.56~{}220.56kBq/mL for KEM) closer to the ground truth (819.79819.79819.79~{}819.79kBq/mL) than OSEM (205.21205.21205.21~{}205.21kBq/mL).
-For the liver lesion in the simulated data, however, use of BSREM-RDPAP actually yielded a smaller activity concentration prediction (261.32261.32261.32~{}261.32kBq/mL) compared to OSEM (267.47267.47267.47~{}267.47kBq/mL).
+These algorithms also yielded higher bone lesion activity concentration predictions than OSEM for both the simulated/clinical data, with the simulated data concentrations (299.11299.11299.11~{}299.11kBq/mL for BSREM-RDPAP and 231.30231.30231.30~{}231.30kBq/mL for KEM) closer to the ground truth (819.79819.79819.79~{}819.79kBq/mL) than OSEM (199.07199.07199.07~{}199.07kBq/mL).
+For the liver lesion in the simulated data, however, use of BSREM-RDPAP actually yielded a smaller activity concentration prediction (244.30244.30244.30~{}244.30kBq/mL) compared to OSEM (251.02251.02251.02~{}251.02kBq/mL).
 This occured because there was no anatomical difference in the surrounding tissue in the liver metastasis, while there was for the bone metastasis.
@@
 The mean predicted counts in the kidneys were 1.764 (STIR) and 1.761 (PyTomography).
-The reconstruction time in PyTomography (5.4 s) was significantly faster than the time required in STIR (51542.2 s), owing to the speed-up factor of GPU implementation.
   2. 2.
@@
 The two images were normalized such that the sum of their voxels was identical.
-1D profiles correspond to the horizontal lines shown on the coronal slices
+1D profiles correspond to the horizontal lines shown on the coronal slices.
+The percent difference in grey and white matter mean uptake between the two software packages were 0.90% and 1.14% respectively.